### v1

# Optimal Stopping with a Predicted Prior

Optimal Stopping with a Predicted Prior
Tian Bai [^1] University of Bergen. Email: tian.bai@uib.no. This work was done when the author was at the University of Hong Kong.    Zhiyi Huang [^2] The University of Hong Kong. Email: zhiyi@cs.hku.hk, dongchen.li@connect.hku.hk    Chui Shan Lee [^3] Hong Kong University of Science and Technology. Email: tracylcs@ust.hk. This work was done when the author was at the University of Hong Kong.    Dongchen Li22footnotemark: 2
(November 2025)
There are two major models of value uncertainty in the optimal stopping literature: the secretary model, which assumes no prior knowledge, and the prophet inequality model, which assumes full information about value distributions. In practice, decision makers often rely on machine-learned priors that may be erroneous. Motivated by this gap, we formulate the model of optimal stopping with a predicted prior to design algorithms that are both consistent, exploiting the prediction when accurate, and robust, retaining worst-case guarantees when it is not.
Existing secretary and prophet inequality algorithms are either pessimistic in consistency or not robust to misprediction. A randomized combination only interpolates their guarantees linearly. We show that a family of bi-criteria algorithms achieves improved consistency-robustness trade-offs, both for maximizing the expected accepted value and for maximizing the probability of accepting the maximum value. We further prove that for the latter objective, no algorithm can simultaneously match the best prophet inequality algorithm in consistency, and the best secretary algorithm in robustness.
1 Introduction
Optimal stopping problems study a fundamental dilemma in decision‑making: to seize an expiring opportunity now, or to wait for a better chance that may never come. A classic example is the secretary problem. In this problem, $n$ values are revealed to a decision maker in a random order. After observing each value, the decision maker immediately decides whether to accept the value. The decision maker can accept at most one value and aims to maximize its expectation. An optimal solution is Dynkin’s algorithm [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989], which rejects the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values, and then accepts the first value that exceeds all previous ones. It guarantees that the expected accepted value is at least $\frac{1}{\mathrm{e}}$ times the maximum. This ratio of the algorithm’s expected accepted value to the maximum value in hindsight is called the competitive ratio.
In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from known prior distributions. In the i.i.d. case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [hillComparisonsStopRule1982, kertzStopRuleSupremum1986, correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021], substantially better than prior-free secretary guarantee of $\frac{1}{\mathrm{e}}\approx 0.368$.
On the one hand, the secretary problem does not capture the abundant data available to decision makers in this era. On the other hand, the assumption of complete knowledge of the prior in prophet inequality is idealized. What real-world decision makers often have are machine-learned and potentially erroneous predicted priors. duttingPostedPricingProphet2019 explored this direction for the general prophet inequality problem. They examined how mispredicted priors affect existing methods and showed a degradation in performance proportional to the prediction error. This can be viewed as an analysis of the smoothness of algorithms under the framework of algorithms with predictions [mitzenmacher2022algorithms].
1.1 Optimal Stopping with a Predicted Prior
We formulate the problem of optimal stopping with a predicted prior and focus on the trade-off between consistency and robustness, i.e., the performance when the predicted prior is correct and when it may be arbitrarily wrong. Consider the optimal stopping problem with $n$ values drawn i.i.d. from an unknown prior distribution $F$. The algorithm is given the number of values $n$ and a predicted prior $\tilde{F}$ in the beginning. Then, the algorithm decides whether to accept each value upon its arrival, based on the history of observed values and the predicted prior $\tilde{F}$. We consider two standard objectives: 1) maximizing the expected accepted value (MaxExp), which we have already discussed, and 2) maximizing the probability of accepting the maximum value (MaxProb).
An algorithm is $\alpha$-consistent if it is $\alpha$-competitive when the predicted prior is correct. It is $\beta$-robust if it remains $\beta$-competitive even when the (mis)predicted prior is arbitrary. Ideally, we want an algorithm to be both $\alpha$-consistent and $\beta$-robust with $\alpha$ and $\beta$ being as large as possible. Such an algorithm leverages prior knowledge when the predicted prior is correct, while retaining meaningful worst-case guarantees.
The $\frac{1}{\mathrm{e}}$-competitive Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust for both objectives. The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by gilbertRecognizingMaximumSequence are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent respectively, but only 0-robust (see Section 2). Further, randomly running either Dynkin’s algorithm or an i.i.d. prophet inequality algorithm linearly interpolates between their consistency and robustness. Finally, the optimality of these competitive ratios implies that no algorithm is better than $\frac{1}{\mathrm{e}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb. These baseline algorithms and impossibilities leave a significant gap between them.
1.2 Overview of Results
This paper introduces two new algorithms for the MaxExp and MaxProb objectives respectively, under a unified framework, and a hardness result for MaxProb. See Figure 1 for an illustration.
(a) MaxExp
(b) MaxProb
Figure 1: Illustration of our results and a comparison with the baselines
Bi-criteria Algorithms.
We define a family of bi-criteria algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Upon observing each value and when no value has been accepted, a bi-criteria algorithm accepts the current value if:
  1. 1.
    It is best-so-far, i.e., greater than or equal to all previous values; (prior-free criterion)
  2. 2.
    Its cdf value w.r.t. $\tilde{F}$ exceeds a time-dependent threshold $\theta(t)$. (prior-dependent criterion)
For now, readers may interpret $t$ as the time step by which the value is observed. Later, we will also consider a continuous-time model to simplify the analysis (see Section 2 for a discussion).
We set the thresholds as follows. Before some time $\lambda_{1}$, the threshold is $1$, and the algorithm rejects all values. After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free best-so-far criterion. These two prior-free phases yield robustness. In fact, we derive this design from first principles, as we will explain shortly in Section 1.3.
We obtain new algorithmic results for both MaxExp and MaxProb under this framework.
Theorem 1 (MaxExp algorithm).
Suppose $0\leq\beta\leq\frac{1}{e}$ and $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. Further, suppose there is a left-continuous non-increasing function $\theta:[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
$\theta(\lambda_{1})$ $~=~1~,$
$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall\lambda_{1}\leq z\leq\lambda_{2}~.$
Then, there is an $\alpha$-consistent $\beta$-robust algorithm for the MaxExp objective.
To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation in Theorem 1, which does not appear to admit a closed-form solution. See Appendix B.
Theorem 2 (MaxProb algorithm).
Suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective for any $\beta\in[0,\frac{1}{e}]$ and
$\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\,\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
We further prove that certain consistency-robustness trade-offs cannot be achieved for the MaxProb objective. In particular, there is no “best-of-both-world” algorithm.
Theorem 3 (MaxProb hardness, informal version of Theorem 15).
There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective only if a polytope $\mathcal{P}(\alpha,\beta)$ is feasible. In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
1.3 Overview of Techniques
From Mispredicted Prior to Misspecified Thresholds.
A key property of bi-criteria algorithms is that they provide two equivalent viewpoints for robustness. By definition, we may consider a real prior $F$ and a (mis)predicted prior $\tilde{F}$ that misleads the algorithm. Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq(F\circ\tilde{F}^{-1}\circ\theta)(t)$. The latter viewpoint is much easier to work with — we just need to restrict the misspecified threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ via our choice of threshold $\theta$. Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ has the same value of 0 or $1$. This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ is between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any value that is best-so-far. See Section 3 for a detailed discussion, and Section 5 for the robustness analysis.
Implicit Sharding.
We show that optimal stopping with a predicted prior is hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest. This is inspired by the sharding technique by harbNewProphetInequalities2025, which breaks each value into many i.i.d. shards based on the prior, such that the original value equals the maximum shard. However, the sharding technique requires full knowledge of the real prior, and thus, does not apply to optimal stopping with a predicted prior. In contrast, implicit sharding directly works with the original values, and only uses the shards implicitly in the analysis through a coupling argument. See Section 4 and specifically Theorem 7 for the formal argument.
Tractable Factor-revealing Linear Programs.
We prove the hardness result by constructing factor-revealing linear programs whose optimal values upper-bound the achievable consistency-robustness trade-offs. Exponential-size linear programs can easily capture all possible algorithms, which map possible sequences of observed values to binary acceptance decisions. The main challenge is to reduce the size without losing much characterization power.
We compress the algorithms’ state space by restricting the real prior to be a conditional distribution of the predicted prior, subject to an upper bound $k$ on the values. Intuitively, the only information an algorithm may conclude from the observed values $x_{1},\ldots,x_{t}$ is $k\geq\max_{1\leq i\leq t}x_{i}$ under this setup. We formalize this intuition, showing that we may, without loss of generality, consider minor-oblivious algorithms whose decisions depend solely on the maximum observed value, not on the detailed observed value sequence. This reduces the state space from an exponential number of possible value sequences to a polynomial number of possible maximum values. See Section 8 for the factor-revealing linear programs and the hardness.
1.4 Related Work
Prophet Inequalities with Full Prior Information.
hillComparisonsStopRule1982 proposed a $(1-\frac{1}{\mathrm{e}})$-competitive algorithm for i.i.d. prophet inequality and the MaxExp objective. This was improved to $0.738$ by abolhassaniBeating11Ordered2017, and finally to approximately $0.745$ by correaPostedPriceMechanisms2017. The latter ratio is tight, matching a hardness result by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986]. For general prophet inequality with non-identical distributions, the optimal competitive ratio is $0.5$ [krengel1978semiamarts, samuel1984comparison]. The prophet secretary problem [esfandiariProphetSecretary2015] interpolates between general and i.i.d. prophet inequalities, with values drawn from non-identical distributions and arriving by a random order. The optimal competitive ratio for prophet secretary is open, with the best lower and upper bounds being $0.688$ [chenProphetSecretaryMatching2025] and $0.723$ [giambartolomeiProphetInequalitiesSeparating2023].
For the MaxProb objective, gilbertRecognizingMaximumSequence gave an optimal algorithm for the i.i.d. case, while samuels1982 derived an explicit expression of the optimal competitive ratio of approximately $0.580$. For non-identical distributions, esfandiariProphetsSecretariesMaximizing2020 obtained an optimal $\frac{1}{\mathrm{e}}$ competitive ratio in the prophet inequality model, and recovered the $0.580$ ratio in the prophet secretary model under a “no-superstar” assumption. nutiSecretaryProblemDistributions2022 later removed this assumption.
See correaRecentDevelopmentsProphet2019 for a survey. We stress that none of these prophet inequality algorithms provides non-trivial robustness.
Prophet Inequalities with Partial Prior Information.
There has been extensive research on the sample-based model, in which the algorithm has access to i.i.d. samples from the prior. One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [azar2014prophet, rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022, ezraProphetInequalitySamples2024]. Notably, the optimal $\frac{1}{2}$-competitive prophet inequality can be achieved with only one sample from each distribution [rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022]. Another line of work studies the sample complexity of learning a nearly optimal algorithm [diakonikolasLearningOnlineAlgorithms2021, correaUnknownIIDProphets2021, correaTwosidedGameGoogol2020, rubinsteinOptimalSinglechoiceProphet2020, cristiProphetInequalitiesRequire2023, guo2021generalizing, correaSampledrivenOptimalStopping2021, jin2024sample], culminating in the state-of-the-art $\tilde{\mathcal{O}}(\frac{n}{\varepsilon})$ sample complexity for i.i.d. prophet inequality [correaSampledrivenOptimalStopping2021], and $\tilde{\mathcal{O}}(\frac{1}{\varepsilon^{2}})$ sample complexity per distribution for the general case [jin2024sample]. cristiProphetInequalitiesRequire2023 gave a unified proof that $\mathcal{O}(\mathrm{poly}(\frac{1}{\epsilon}))$ samples per distribution suffice for several major variants of prophet inequalities.
Other models of partial information include learning nearly optimal algorithms via quantile queries [liProphetInequalityIID2023], bandit feedback [gatmiry2024bandit], or simple posted-pricing policies [perezsalazarIIDProphetInequality2023]. These models and the sample-based model are incomparable to ours. They assume partial information from the correct prior, while the model of this paper considers a predicted prior that may be arbitrarily wrong.
Algorithms with Predictions.
Our work falls within the algorithms with predictions framework [mitzenmacher2022algorithms] for augmenting worst-case analysis of algorithms with machine-learned advice. Consistency, robustness, and smoothness are the primary performance metrics, capturing the algorithms’ performance when the prediction is correct, completely wrong, and partially wrong. The framework has been applied to many optimization problems, including scheduling [kumar2018improving] and caching [lykouris2021competitive]. In the context of optimal stopping, antoniadisSecretaryOnlineMatching2020 studied a prediction of the maximum value, fujiiSecretaryProblemPredictions2024 assumed having a predicted sequence of values, and braun2024secretary considered a prediction of the gap between the maximum value and the $k$-th largest. In contrast, this paper considers distributional predictions, which aligns with the literature of prophet inequality and the machine-learned priors in practice. As mentioned earlier, duttingPostedPricingProphet2019 also considered mispredicted prior with a focus on smoothness, while this paper examines the trade-off between consistency and robustness.
2 Preliminaries
Let $\mathbb{N}$ denote the set of positive integers. For any $n\in\mathbb{N}$, define $n]\coloneqq\{1,2,\ldots,n\}$. Let $\mathcal{D}$ be the set of all probability distributions supported on $0,\infty)$. We adopt the standard abbreviations cdf, pdf, and pmf for cumulative distribution function, probability density function, and probability mass function. We often use the same symbol for a distribution and its cdf (e.g., $F$), when clear from context. Unless stated otherwise, expectations and probabilities are taken over the randomness of the values and any internal randomness of the algorithm.
2.1 Optimal Stopping
Consider $n$ values $x_{1},x_{2},\ldots,x_{n}$ drawn i.i.d. from a prior distribution $F\in\mathcal{D}$. The values arrive one by one. Upon the arrival of each value $x_{i}$, the algorithm observes the realized value, and then must immediately decide whether to accept $x_{i}$ and stop, or to reject it and continue. At most one value may be accepted.
Fix any algorithm, let Alg be the index of the value it accepts, and thus, $x_{\textsc{Alg}}$ be the corresponding accepted value. We define $x_{\textsc{Alg}}=0$ if the algorithm rejects all values.
We consider the maximization problem, where larger values are preferable. The literature has studied two objectives. Let $x_{\textsc{Opt}}$ denote the maximum value among all $n$ values, i.e., $x_{\textsc{Opt}}\coloneqq\max_{i\in[n]}x_{i}$. The max expectation (MaxExp) objective considers the expected value of the algorithm’s accepted decision, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$. The corresponding competitive ratio is
$\frac{\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]}{\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]}~.$
The max probability (MaxProb) objective considers the probability of accepting the maximum:
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~.$
This coincides with the competitive ratio since the offline optimal choice is always maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
2.2 Optimal Stopping with a Predicted Prior
We assume that the number of values $n$ is known. Depending on what the algorithm knows about the prior distribution $F$, we have different optimal stopping problems. If nothing is known, we have essentially the secretary problem [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989]. [^4] The classical secretary problem considers random order without prior. Nonetheless, correaProphetInequalitiesIID2019 showed that the optimal competitive ratio is still $\frac{1}{\mathrm{e}}$ for an unknown prior. With complete information of $F$, we have the i.i.d. prophet inequality [correaUnknownIIDProphets2021].
This paper considers a new model between the two extremes, which we dub optimal stopping with a predicted prior. In this model, the algorithm is given a possibly erroneous predicted prior $\tilde{F}$. Hence, upon receiving each value $x_{i}$, the algorithm decides whether to accept based on the number of values $n$, the predicted prior $\tilde{F}$, and the realizations of the values observed thus far including $x_{i}$.
Following the literature of algorithms with predictions [mitzenmacher2022algorithms], we consider the consistency and robustness of stopping algorithms.
Definition 1 (Consistency).
An algorithm is $\alpha$-consistent if for any prior distribution $F\in\mathcal{D}$, it is $\alpha$-competitive when the predicted prior is accurate, i.e., when $\tilde{F}=F$.
Definition 2 (Robustness).
An algorithm is $\beta$-robust if for any prior distribution $F\in\mathcal{D}$, it is $\beta$-competitive regardless of which predicted prior $\tilde{F}$ is given.
2.3 Simplifying Assumptions for Algorithm Analyses
For our algorithmic results, we adopt two standard assumptions from the literature that simplify the analysis and are without loss of generality.
Continuous-time Random-order Model.
Besides the above classical model, which we will refer to as the discrete-time model, we may artificially introduce a random order of arrivals by letting each value $x_{i}$ with $i\in[n]$ arrive in a continuous time horizon $[0,1]$, drawing its arrival time $t_{i}$ independently and uniformly from $[0,1]$. Since the values are drawn i.i.d. from $F$, the problem remains the same.
An algorithm for the discrete-time model also works in the continuous-time model by ignoring the arrival time. The converse is also true. We can simulate the continuous-time model by first drawing $n$ arrival times independently and uniformly from $[0,1]$, and then mapping these arrival times in ascending order to the values in their arrival order.
Full-support Continuous Priors.
For the analysis, we further assume that both $F$ and $\tilde{F}$ are continuous with full support on $0,\infty)$. This assumption is without loss of generality, e.g., by independently perturbing each value with non-negative noise drawn from an exponential distribution with parameter $\varepsilon\to 0$; see correaProphetSecretaryBlind2021 for a similar treatment.
For any full-support continuous distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
$F^{-1}(q)\coloneqq\inf\big\{x\geq 0:F(x)\geq q\big\},\quad\forall q\in[0,1].$
Importantly, $F$ and $F^{-1}$ are strictly increasing on the interiors of their domains. Further, we have the boundary conditions $F^{-1}(0)=0$ and $F(0)=0$, and by convention $F^{-1}(1)=\infty$ and $F(\infty)=1$.
3 Bi-criteria Algorithms
We start with two definitions that will be repeatedly used throughout the paper.
Definition 3.
The prefix maximum for any time $t\in[0,1]$ is the maximum value arrived before time $t$, i.e., $\max_{i\,:\,t_{i}<t}x_{i}$, with the convention that it equals 0 if no value arrives before $t$.
Definition 4.
A value $x$ arriving at time $t$ is best-so-far if $x$ is greater than or equal to the prefix maximum for time $t$.
We consider the following family of algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Throughout the paper, threshold functions are always non-increasing maps from $[0,1]$ to $[0,1]$.
Bi-criteria Algorithm $\mathcal{A}(\tilde{F},\theta)$ For each value $x_{i}$ arriving at time $t_{i}$ when no value has been accepted, accept $x_{i}$ if: 1. Value $x_{i}$ is best-so-far; and 2. The cdf value of $x_{i}$ w.r.t. $\tilde{F}$ is greater than the threshold at time $t_{i}$, i.e., if $\tilde{F}(x_{i})>\theta(t_{i})$.
The first prior-free criterion is without loss of optimality for MaxProb and is critical for the robustness guarantees for both MaxProb and MaxExp. The second prior-dependent criterion represents how the algorithm uses the predicted prior to guide its decisions.
Specifically, if $\theta(t)=1$, the algorithm accepts values with probability 0 at time $t$, regardless of how large the values are. If $\theta(t)=0$, the algorithm accepts any value that is best-so-far at time $t$, regardless of the value’s relative standing w.r.t. the predicted prior.
Algorithms under this framework satisfy many properties, including a simple necessary and sufficient condition for the algorithm to reject all values before time $t$, as follows. As discussed in Section 2.3, we assume the distribution is continuous and thus ties occur with probability zero.
Proposition 4.
The algorithm $\mathcal{A}$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
Proof.
If $\tilde{F}(y)\leq\theta(s)$, then all values $x$ arriving before time $s$ are rejected because the cdf value is below the threshold, as $\tilde{F}(x)\leq\tilde{F}(y)\leq\theta(s)\leq\theta(t)$ for any $t\leq s$. Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be best-so-far.
If $\tilde{F}(y)>\theta(s)$, then $y$ is best-so-far and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
Corollary 5.
Suppose a value $x$ arrives at time $t$, and its prefix maximum $y$ arrives at time $s<t$. The algorithm accepts $x$ if and only if 1) $\tilde{F}(x)>\theta(t)$, 2) $x\geq y$, and 3) $\tilde{F}(y)\leq\theta(s)$.
3.1 Examples
This family encompasses many classical algorithms as special cases. We present three examples.
Example 1 (Dynkin’s algorithm [dynkinOptimumChoiceInstant1963], continuous-time variant).
Dynkin’s algorithm accepts the first best-so-far value after rejecting the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values. The continuous-time variant (e.g., feldmanImprovedCompetitiveRatios2011) accepts the first best-so-far value after time $\frac{1}{\mathrm{e}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\frac{1}{\mathrm{e}}]$, and $\theta(t)=0$ for $t\in(\frac{1}{\mathrm{e}},1]$.
Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-competitive for both MaxProb and MaxExp, without any information about the prior [dynkinOptimumChoiceInstant1963, feldmanImprovedCompetitiveRatios2011]. Hence, it is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
Example 2 (Gilbert-Mosteller algorithm [gilbertRecognizingMaximumSequence], continuous-time variant).
Gilbert and Mosteller calculated the optimal thresholds for MaxProb through backward induction with a correct prior, i.e., $\tilde{F}=F$. Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to samuels1982: [^5] The competitive ratio presented in samuels1982 is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
$\int_{0}^{1}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
The continuous-time variant under our framework corresponds to the threshold function given by the following equation:
$\int_{t}^{1}\frac{(1-s+s\theta(t))^{n-1}-\theta(t)^{n-1}}{1-s}\,\mathrm{d}s=\theta(t)^{n-1}~.$
It achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm (as a special case of Theorem 2). However, the algorithm is merely 0-robust. To illustrate, consider the case where the prior $F$ is uniform over $[0,1]$, whereas the predicted prior $\tilde{F}$ is uniform over $[2,3]$. In this case, the algorithm rejects all values, as they have cdf value 0 w.r.t. the predicted prior.
Example 3 (Single-threshold algorithm [ehsaniProphetSecretaryCombinatorial2018], i.i.d. case).
Let $\theta(t)=1-\tfrac{1}{n}$ for $t\in[0,1]$. [^6] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\frac{1}{\mathrm{e}}$, which corresponds to $F(x_{i})\approx 1-\frac{1}{n}$ for each $i$ in the i.i.d. case.
The single-threshold algorithm achieves a competitive ratio of $1-\frac{1}{\mathrm{e}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp. Hence, it is $(1-\frac{1}{\mathrm{e}})$-consistent for MaxExp. However, it suffers from the same counterexample and is only 0-robust like Gilbert-Mosteller algorithm.
In sum, the threshold functions corresponding to existing algorithms fail to achieve nontrivial trade-offs between consistency and robustness. They either ignore the predicted prior and have pessimistic consistency (Dynkin’s algorithm) or rely on having an accurate prior without any robustness (Gilbert-Mosteller algorithm and single-threshold algorithm).
3.2 Robust Threshold Functions
We next explain our design of robust threshold functions. First, we need to understand how an adversary would choose a worst-case prior distribution $F$, given the predicted prior $\tilde{F}$, threshold function $\theta$, and the corresponding algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$. The key observation is the following lemma.
Lemma 6 (From mispredicted prior to misspecified thresholds).
For any distributions $F$ and $\tilde{F}$, and any threshold function $\theta$, algorithms $\mathcal{A}\big(\tilde{F},\theta\big)$ and $\mathcal{A}(F,F\circ\tilde{F}^{-1}\circ\theta)$ are the same.
Proof.
The lemma holds because $\tilde{F}(x)\,{>}\,\theta(t)$ if and only if $F(x)\,{>}\,F\circ\tilde{F}^{-1}\circ\theta(t)$, according to the monotonicity of $F$ and $\tilde{F}^{-1}$ and their boundary conditions. ∎
Despite its simple one-sentence proof, the lemma has an important implication. Instead of considering algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ parameterized with a mispredicted prior $\tilde{F}$, we may equivalently consider algorithm $\mathcal{A}\big(F,\theta\big)$ parameterized with the correct prior $F$ and a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$.
Importantly, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$ is not arbitrary. By the boundary conditions of $F$ and $\tilde{F}^{-1}$, if the original threshold $\theta(t)$ equals 0 or $1$ at time $t$, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta(t)$ must have the same value 0 or $1$. Further, it is easy to show that this is the only constraint. Let $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$. Then, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta(t)$ can be any non-increasing function for $t$ between $\lambda_{1}$ and $\lambda_{2}$, while having value $1$ for $0\leq t<\lambda_{1}$ and value 0 for $\lambda_{2}<t\leq 1$.
We will further show in Section 5 that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$. They correspond to a variant of Dynkin’s algorithm that starts to accept best-so-far values at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\frac{1}{\mathrm{e}}$. The competitive ratio $\lambda\ln\lambda$ of this algorithm (e.g., feldmanImprovedCompetitiveRatios2011) motivates the following definition of robust threshold functions.
Definition 5.
Consider any feasible robustness parameter $0\leq\beta\leq\frac{1}{\mathrm{e}}$. Let $0\leq\lambda_{1}\leq\frac{1}{\mathrm{e}}\leq\lambda_{2}\leq 1$ be the solutions to:
$\lambda\ln\lambda~=~\beta~.$
Readers familiar with the Lambert $\mathcal{W}$ function may note that:
$\lambda_{1}\coloneqq\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}~,\quad\qquad\lambda_{2}\coloneqq\mathrm{e}^{\mathcal{W}_{0}(-\beta)}~.$ (1)
For any threshold function $\theta$, define a corresponding robust threshold function as:
$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t<\lambda_{2}~;\\ 0&\lambda_{2}\leq t\leq 1~.\end{cases}$
For example, Figure 2 shows the robust counterpart of the threshold function from Gilbert-Mosteller algorithm.
Figure 2: Illustration of the robust threshold function corresponding to Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{0}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.538$.
3.3 Brief Discussion on Algorithms Outside the Framework
In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is best-so-far. With a correct prior, the optimal policy by backward induction takes this form, where the threshold is the expected accepted value if the algorithm rejects the current value. For example, see the work by krengel1977semiamarts on general prophet inequality and the more recent research [correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021, harbNewProphetInequalities2025] for i.i.d. prophet inequality. These algorithms are outside our framework, where the taken value must be best-so-far, in addition to surpassing the time-dependent threshold.
We stress that the best-so-far requirement is critical for robustness. The mentioned algorithms that have no best-so-far requirements are only 0-robust. To see this, consider a predicted prior supported on values much smaller than those in the real prior. Then, all values from the real prior will have cdf value essentially $1$ w.r.t. the predicted prior. As a result, the algorithm accepts the first value arrived when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
Meanwhile, our framework is limited in that it does not adjust thresholds adaptively based on the observed values. Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [dynkinOptimumChoiceInstant1963] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.5801$-competitive algorithm [gilbertRecognizingMaximumSequence] in the MaxProb setting or the $0.745$-competitive algorithm [correaProphetInequalitiesIID2019] in the MaxExp setting if the observed values are consistent with the predicted prior. We consider a static threshold function to make the analyses tractable. Exploring algorithms with adaptive thresholds is an interesting direction for future work.
4 Implicit Sharding
This section introduces an implicit sharding technique that reduces instances with a smaller number of values to instances with arbitrarily many values. In this sense, it suffices to design algorithms for the limit case when the number of values $n$ tends to infinity. The main result of this section is the following theorem.
Theorem 7.
For any $n,k\in\mathbb{N}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
By our convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$. Importantly, the maximum of $k$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$ follows distribution $\tilde{F}$. We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$. Let $x_{1},\dots,x_{n}$ be the $n$ values by their arrival order in the original instance. Further, consider an instance of $nk$ values $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ drawn i.i.d. from distribution $F^{\nicefrac{{1}}{{k}}}$ with sorted arrival times $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ drawn i.i.d. and uniformly from $[0,1]$.
Definition 6 (Coupling for Sharding).
Let $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$, and let $s_{i}$ be the arrival time of $\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}$. We use $s_{i}$ instead of $t_{i}$ to avoid confusion, as $s_{i}$ is not uniform over $[0,1]$.
Naïvely, we would like to simulate algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ by sampling $x^{\prime}_{ij}$, $j\in[k]$, from $F^{\nicefrac{{1}}{{k}}}$ conditioned on the realization of $x_{i}$, and accept value $x_{i}$ if and only if one of the $x^{\prime}_{ij}$’s is accepted by $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. This is the sharding technique by harbNewProphetInequalities2025 for the prophet inequality setting. Unfortunately, the algorithm does not know $F$ in the problem of optimal stopping with a predicted prior, and thus cannot sample $x^{\prime}_{ij}$ from $F^{\nicefrac{{1}}{{k}}}$. Sampling $x^{\prime}_{ij}$ based on the predicted prior $\tilde{F}$ and the corresponding $\tilde{F}^{\nicefrac{{1}}{{k}}}$ preserves the consistency of $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, but not robustness.
We propose an implicit sharding method that partially simulates algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ only at the arrival times of the $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$. We will show that this partial simulation is at least as good as the full simulation in the original sharding method. Further, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis.
Bi-criteria Algorithm with Implicit Sharding $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ 1. Sample nknk arrival times i.i.d. from $\mathrm{Unif}[0,1]$ and sort them as:
             $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ .
2. For each $i\in[n]$, sample $s_{i}$ uniformly from $\{t^{\prime}_{i1},t^{\prime}_{i2},\ldots,t^{\prime}_{ik}\}$. 3. For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if: (a) Value $x_{i}$ is best-so-far; and (b) The cdf value of $x_{i}$ w.r.t. $\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if:
              $\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})~.$
Proof of Theorem 7.
It suffices to show that for any real prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
Suppose algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ accepts value $x_{i}$, and $x^{\prime}_{ij}=x_{i}$ is the corresponding value in the coupled instance for algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. Since $x_{i}$ is best-so-far, $x^{\prime}_{ij}$ is also best-so-far because
$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}\leq i,\;j^{\prime}\in[k].$
Further, the cdf value of $x^{\prime}_{ij}$ is greater than the threshold at time $t^{\prime}_{ij}$, as
$\tilde{F}^{\nicefrac{{1}}{{k}}}(x^{\prime}_{ij})=\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})=\theta(t^{\prime}_{ij})~.$
Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is best-so-far. Taking expectations over the coupling completes the proof. ∎
5 Proof of Robustness
We analyze the robustness of our algorithm for MaxProb (which then implies robustness for MaxExp) in a stronger setting that fixes the realized values and randomizes only the arrivals. This problem is also known as the game of Googol [fergusonWhoSolvedSecretary1989, gnedin1994solution].
Proposition 8.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value with probability at least
$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}~,$
over the random arrivals of the values.
This implies the robustness guarantees for both MaxProb and MaxExp as corollaries through a standard argument.
Corollary 9.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, and any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp.
Proof.
By the law of total probability, we have
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},\dots,x_{n}}\Big[\,\operatorname{\mathbf{Pr}}_{t_{1},\dots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},\dots,x_{n}\,\big]\,\Big]~.$
By Proposition 8, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity. This proves robustness for MaxProb. Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, i.e., the same robustness for MaxExp. ∎
Therefore, it remains to prove Proposition 8. We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value. Without loss of generality (by relabeling the distinct realized values), assume $x_{1}<x_{2}<\dots<x_{n}$, so $x_{\textsc{Opt}}=x_{n}$.
Lemma 10.
For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
$\int_{0}^{1}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{n})>\theta(t)\}}}\bigg((1-t)^{n-1}+\sum_{i=1}^{n-1}(1-t)^{n-1-i}\int_{0}^{t}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(s)\}}}\,\mathrm{d}s\bigg)\,\mathrm{d}t~.$ (2)
Proof.
Since $x_{n}$ is always best-so-far, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$. By Proposition 4, the latter is equivalent to the event that the prefix maximum for time $t_{n}$ has cdf under $\tilde{F}$ at most the threshold at its arrival time. Hence, the acceptance probability is
$\int_{0}^{1}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{n})>\theta(t_{n})\}}}\operatorname{\mathbf{Pr}}\big[\,\text{prefix maximum for time $t_{n}$ is below threshold}\,\big]\,\mathrm{d}t_{n}~.$
To compute the expression for the probability in the integrand, consider each possible prefix maximum separately. First, the prefix maximum is 0 if no value arrives before $t_{n}$, which occurs with probability $(1-t_{n})^{n-1}$. Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$. In this case, we also require $\tilde{F}(x_{i})\leq\theta(t_{i})$ for the arrival time $t_{i}\in[0,t_{n})$, contributing
$(1-t_{n})^{n-1-i}\int_{0}^{t_{n}}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(t_{i})\}}}\,\mathrm{d}t_{i}~.$
Putting together and changing variables $t_{n}$ with $t$ and $t_{i}$ with $s$, the lemma follows. ∎
We are now ready to prove Proposition 8.
Proof of Proposition 8.
By Lemma 10, it suffices to show that the expression in Equation 2 is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for any predicted prior $\tilde{F}$ and threshold function $\theta$.
First, we fix $\theta$ and characterize the worst-case predicted prior $\tilde{F}$. Equation 2 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\dots\leq\tilde{F}(x_{n})\leq 1$. Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n-1})$. Therefore, given any fixed threshold function $\theta$, Equation 2 is minimized when all $\tilde{F}(x_{i})$ are equal to some $q\in[0,1]$.
In this case, the threshold function is effectively binary. A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n})$ all surpass the threshold. Similarly, a threshold $\theta(t)\geq q$ is equivalent to $\theta(t)=1$, as none surpasses it. Hence, we may without loss of generality consider a binary threshold, such that for some $\lambda\in[\lambda_{1},\lambda_{2}]$, it holds
$\theta(t)=\begin{cases}1,&\text{if }t\leq\lambda~;\\ 0,&\text{otherwise~.}\end{cases}$
This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\frac{1}{\mathrm{e}}$. This is $(-\lambda\ln\lambda)$-competitive (e.g., feldmanImprovedCompetitiveRatios2011). Since $\lambda\ln\lambda$ is concave in $\lambda$ on $[0,1]$, its minimum over $\lambda_{1},\lambda_{2}]$ is attained at an endpoint, yielding the bound $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$. ∎
6 Algorithm for MaxExp: Proof of Theorem 1
By the implicit sharding method (Theorem 7), it suffices to design an algorithm $\mathcal{A}(\tilde{F},\theta)$ under the bi-criteria framework, and prove the desired consistency and robustness guarantees as the number of values $n$ tends to infinity. Further, by Corollary 9, the robustness guarantee holds for any robust threshold $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$.
The rest of the section focuses on the consistency guarantee, for any fixed robustness parameter $\beta\in[0,\frac{1}{e}]$ and the corresponding $\lambda_{1},\lambda_{2}$. Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will only use $F$ throughout the section. To formally study the limit as $n\to\infty$, we will design the threshold function for the extremal distribution $F^{n}$. More precisely, we will design a threshold function $\theta$ and analyze the consistency of algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ for $n$ values, and in particular, the limit when $n$ tends to infinity.
As a standard approach, we design the algorithm so that for any $\ell\geq 0$, it holds
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~.$ (3)
This implies the consistency guarantee as
$\operatorname{\mathbf{E}}\big[\,x_{\textsc{Alg}}\,\big]~=~\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]\,\mathrm{d}\ell~\geq~\alpha\cdot\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]\,\mathrm{d}\ell~=~\alpha\cdot\operatorname{\mathbf{E}}\big[\,x_{\textsc{Opt}}\,\big]~.$
6.1 Sufficient Conditions for Infinitely Many Values
We consider a general threshold function $\theta$ and derive an expression for the left-hand side of Equation 3, i.e., the probability that the algorithm accepts a value that is at least $\ell$.
Lemma 11.
For any threshold function $\theta$, any number of values $n$, and any $\ell\geq 0$, algorithm $\mathcal{A}\big(F,\theta^{\nicefrac{{1}}{{n}}}\big)$ satisfies that
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~=~\int_{F(\ell)^{n}}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\Big(1-t+t\cdot\min\big\{\theta(s),q\big\}^{\nicefrac{{1}}{{n}}}\Big)^{n-1}q^{-\frac{n-1}{n}}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~,$
where $\theta^{-1}(x)\coloneqq\inf\{t\in[0,1]\colon\theta(t)<x\}$ is the generalized inverse of $\theta$, defined as the boundary above which $\theta(\cdot)$ is strictly smaller than $x$.
Proof.
Consider any value $x_{i}$ arriving at time $t_{i}$, with the prefix maximum for time $t_{i}$ being $y$ and arriving at time $s$. By Corollary 5, the algorithm $\mathcal{A}\big(F,\theta^{\nicefrac{{1}}{{n}}}\big)$ accepts $x_{i}$ if and only if
$F(x_{i})>\theta(t_{i})^{\nicefrac{{1}}{{n}}}~,\quad x_{i}\geq y~,\quad F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}~.$
We consider the probability that $x_{i}$ is at least $\ell$ and accepted by the algorithm, enumerating over all triples $(x_{i},t_{i},s)$. First, value $x_{i}$ follows distribution $F$ and satisfies $x_{i}\geq\ell$. Further, its arrival time $t_{i}$ is uniform over $[0,1]$, and satisfies $F(x_{i})>\theta(t_{i})^{\nicefrac{{1}}{{n}}}$, or equivalently, $F(x_{i})^{n}>\theta(t_{i})$. Finally, given $x_{i}$ and $t_{i}$, the arrival time $s$ of the prefix maximum is uniformly distributed over $0,t_{i}]$. We may artificially sample $s$ from this distribution even if no values arrive before $t_{i}$. In sum, value $x_{i}$ is at least $\ell$ and accepted by the algorithm with probability
$\int_{x_{i}=\ell}^{x_{i}=\infty}\int_{\theta^{-1}(F(x_{i})^{n})}^{1}\int_{0}^{t_{i}}\frac{1}{t_{i}}\,\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}\mid x_{i},t_{i},s\,\big]\,\mathrm{d}s\,\mathrm{d}t_{i}\,\mathrm{d}F(x_{i})~.$
For each of the n1n-1 values other than $x_{i}$, either it arrives after time $t_{i}$, in which case there is no constraint on the realized value, or it arrives before $t_{i}$, in which case the cdf of the realized value is at most $\min\{\theta(s)^{\nicefrac{{1}}{{n}}},F(x_{i})\}$ by the upper bounds for $y$. Thus,
$\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}\mid x_{i},t_{i},s\,\big]~=~\big(\,1-t_{i}+t_{i}\min\{\theta(s)^{\nicefrac{{1}}{{n}}},F(x_{i})\}\,\big)^{n-1}~.$
Putting together and changing variables with $t=t_{i}$ and $q=F(x_{i})^{n}$, the above probability is
$\frac{1}{n}\int_{F(\ell)^{n}}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\Big(1-t+t\cdot\min\big\{\theta(s),q\big\}^{\nicefrac{{1}}{{n}}}\Big)^{n-1}q^{-\frac{n-1}{n}}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~.$
Multiplying by $n$ by symmetry across all $x_{i}$’s proves the lemma. ∎
For any $y\in[0,1]$, if we let $n$ tend to infinity while keeping the same $F(\ell)^{n}=y$, the desired Inequality (3) becomes
$\int_{y}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
where we use the fact that $\big(1-t+tz^{\nicefrac{{1}}{{n}}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity. [^7] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n$ tends to infinity and consider any $\varepsilon>0$. By the uniform convergence of the stated limit for $q\in[\varepsilon,1]$, we can apply implicit sharding to ensure Equation 3 for $\alpha=\alpha^{*}(1-\varepsilon)$ and any $q=F(\ell)^{n}\in[\varepsilon,1]$. Finally, Equation 3 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most a $1-\varepsilon$ factor.
Define the function with support set $[0,1]$ as
$g(q)~\coloneqq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
Then, it suffices to design a threshold function $\theta$ such that for any $q\in[0,1]$, it holds
$g(q)~\geq~\alpha~.$
6.2 Proof of Theorem 1
Next, we focus on a robust threshold function $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, and for ease of notation we continue to write $\theta$ instead of $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, restricting $\theta$ from now on such that
  1. 1)
    $\theta(t)=1$ for $t\in[0,\lambda_{1})$; and
  2. 2)
    $\theta(t)=0$ for $t\in(\lambda_{2},1]$.
Further, recall the assumed properties in Theorem 1:
  1. 3)
    $\theta$ is left-continuous, which implies $\theta(\lambda_{1})=1$, and non-increasing; and
  2. 4)
    For any $z\in[\lambda_{1},\lambda_{2}]$, we have
    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\theta(z)~.$
Consider any $q$ with $\theta(\lambda_{2})\leq q\leq\theta(\lambda_{1})=1$. Let $z=\theta^{-1}(q)$, and then by property 4), we have
$\alpha$ $~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\theta(\max\{s,\theta^{-1}(q)\})^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
$~=~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),\theta(\theta^{-1}(q))\big\}^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t.$
By left-continuity of $\theta$ and the definition of $\theta^{-1}$, we have $\theta(\theta^{-1}(q))\geq q$. Therefore,
$\alpha~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\theta(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q).$
Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\theta(\lambda_{2}))$ by reducing it to the first case.
Lemma 12.
For any $q\in[0,\theta(\lambda_{2})]$, we have
$g(q)\geq g\big(\theta(\lambda_{2})\big).$
Proof.
For any $q$ in the range $0,\theta(\lambda_{2})]$, we have $\theta^{-1}(q)=\lambda_{2}$. Also, observe that $\theta(s)\geq\theta(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\theta(s)=0$ for $s\in[\lambda_{2},1]$. Hence,
$g(q)=\int_{\lambda_{2}}^{1}\int_{0}^{\lambda_{2}}\frac{1}{t}q^{-(1-t)}\,\mathrm{d}s\,\mathrm{d}t=\int_{\lambda_{2}}^{1}\frac{\lambda_{2}}{t}q^{-(1-t)}\,\mathrm{d}t~,$
which is decreasing in $q$. ∎
7 Algorithm for MaxProb: Proof of Theorem 2
We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value when the predicted prior is correct.
Lemma 13.
For any $n\in\mathbb{N}^{+}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ is
$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s$
competitive for MaxProb.
Proof.
Consider any value $x_{i}$ and the probability that $x_{i}$ is both maximum and accepted by the algorithm. Recall that $t_{i}$ is the arrival time of $x_{i}$, and let $s$ denote the arrival time of the prefix maximum $y$ for time $t_{i}$. By Corollary 5, the algorithm accepts $x_{i}$ if and only if
$F(x_{i})>\theta(t_{i})~,\quad x_{i}\geq y~,~\text{and}\quad F(y)\leq\theta(s)~.$
We enumerate over all possible triples $(x_{i},t_{i},s)$. First, arrival time $t_{i}$ distributes uniformly over $[0,1]$. Further, value $x_{i}$ follows distribution $F$, and must satisfy $F(x_{i})\geq\theta(t_{i})$ for $x_{i}$ to be accepted. Finally, given $x_{i}$ and $t_{i}$, the arrival time $s$ of the prefix maximum is uniformly distributed over $[0,t]$. We may artificially sample $s$ from this distribution even if no values arrive before $t_{i}$. Therefore, the probability that $x_{i}$ is both maximum and accepted equals
$\int_{0}^{1}\int_{\theta(t_{i})}^{1}\int_{0}^{t_{i}}\frac{1}{t_{i}}\,\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)\,\,,\,x_{\textsc{Opt}}=x_{i}\mid x_{i},t_{i},s\,\big]\,\mathrm{d}s\,\mathrm{d}F(x_{i})\,\mathrm{d}t_{i}~.$
For each of the n1n-1 values other than $x_{i}$, either it arrives before time $t_{i}$, in which case its cdf value must be at most $\min\{\theta(s),F(x_{i})\}$ by the upper bounds for the prefix maximum $y$; or it arrives after $t_{i}$, in which case it must be at most $x_{i}$ since $x_{i}$ is maximum. Therefore,
$\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)\,\,,\,x_{\textsc{Opt}}=x_{i}\mid x_{i},t_{i},s\,\big]~=~\Big(\,t_{i}\min\big\{\theta(s),F(x_{i})\big\}+(1-t_{i})F(x_{i})\,\Big)^{n-1}~.$
Putting together, changing variables with $t=t_{i}$ and $q=F(x_{i})$, and multiplying by $n$ by symmetry across all $x_{i}$’s, the algorithm accepts the maximum value with probability
$n\int_{0}^{1}\int_{\theta(t)}^{1}\int_{0}^{t}\frac{1}{t}\,\big(\,t\min\{\theta(s),q\}+(1-t)q\,\big)^{n-1}\,\mathrm{d}s\,\mathrm{d}q\,\mathrm{d}t~.$
Changing the order of integration and simplifying yields the stated expression for $\Gamma_{n}(\theta)$. ∎
Lemma 14.
For any $n\in\mathbb{N}^{+}$ and any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, define threshold function $\theta_{n}^{*}$ such that for any $s\in[0,1]$, $\theta_{n}^{*}(s)$ is the unique solution to
$\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t~=~\theta_{n}^{*}(s)^{n-1}~.$ (4)
We have
$\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\geq~-\lambda_{1}\ln\lambda_{1}+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where equality holds when $n$ tends to infinity.
Note that $\theta_{n}^{*}$ is exactly the threshold function used in the continuous-time Gilbert-Mosteller algorithm (Example 2). We defer the proofs of some lemmas to Appendix A.
Proof.
By the expression for $\Gamma_{n}(\theta)$ in Lemma 13, we derive the definition of $\theta_{n}^{*}$ from the first-order condition for $\theta(t)$, i.e.,
[…]
Further, $\theta_{n}^{*}(s)$ is decreasing in $s$ because Equation 4 can be rewritten as
$\int_{s}^{1}\frac{\big(\tfrac{1-t}{\theta_{n}^{*}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t~=~1~,$ (5)
except for the boundary condition $\theta_{n}^{*}(1)=0$.
In sum, $\theta_{n}^{*}$ is a threshold function satisfying the first-order conditions for $t\in[0,1]$. Therefore, $\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is a threshold function satisfying the first-order conditions for $t\in[\lambda_{1},\lambda_{2}]$ on one hand, and the restrictions of being $1$ for $t\in[0,\lambda_{1})$ and 0 for $t\in(\lambda_{2},1]$ on the other hand.
We next write the competitive ratio $\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ as
$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{*}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
We now focus on the limit case when $n$ tends to infinity. This is without loss of generality, either by the implicit sharding technique, or by the fact that the above ratio is non-increasing in $n$ (see Lemma 19 in Section A.1). In particular, the latter means that the stated consistency and robustness guarantees also hold directly for the bi-criteria algorithm without implicit sharding.
When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}$ by basic calculus, and the last term tends to 0. It remains to show that, as $n\to\infty$, the integrand of the second term converges to
$\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t~.$
By the definition of $\theta_{n}^{*}$ (Equation 4), the integrand equals
$\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{*}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t$
$\qquad={}\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n-1}}{t}\,\mathrm{d}t~.$
Finally, we prove in Section A.2 (Lemma 20) that
$\theta_{n}^{*}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$
uniformly on compact subsets of $[0,1)$, which yields the above limit as $n\to\infty$. [^8] Rigorously, for any $\varepsilon>0$, uniform convergence over $0,1-\varepsilon]$ implies the convergence of $\int_{s}^{1-\varepsilon}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t$. The difference from the ideal term is bounded by $\int_{1-\varepsilon}^{1}\frac{1}{t}\,\mathrm{d}t=\mathcal{O}(\varepsilon)$, which vanishes as $\varepsilon\to 0$.
8 Hardness for MaxProb
This section proves that certain consistency-robustness trade-offs for MaxProb are unattainable. We consider the discrete-time model and discrete distributions. In the presence of ties, we adopt the all-ties-win convention, i.e., the algorithm only needs to accept any value that equals $x_{\textsc{Opt}}$. This convention only strengthens the hardness result.
Theorem 15.
If there is an $\alpha$-consistent $\beta$-robust algorithm for MaxProb, then for any integers $n,K\in\mathbb{N}$ and any predicted prior $\tilde{F}$ over $[K]$, the following polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$ for variables $\{y_{t,\ell}\}_{t\in[n],\,\ell\in[K]}$ is feasible:
$\forall\ell,\quad$ $y_{0,\ell}=1;$ (6)
$\forall t,\forall\ell,\quad$ $0\leq y_{t,\ell}\leq 1;$ (7)
$\forall t,\forall\ell,\quad$ $\Delta^{t}(\ell)\,y_{t,\ell}-\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}\geq 0$ (8)
$\forall t,\forall\ell,\quad$ $\Delta^{t}(\ell)\,y_{t,\ell}-\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}\leq\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m};$ (9)
$\sum_{t\in[n]}\sum_{\ell\in[K]}\tilde{F}(\ell)^{n-t}\Bigg(\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m}+\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}-\Delta^{t}(\ell)\,y_{t,\ell}\Bigg)\geq\alpha;$ (10)
$\forall k,\quad$ $\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\Bigg(\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m}+\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}-\Delta^{t}(\ell)\,y_{t,\ell}\Bigg)\geq\beta,$ (11)
where $\tilde{f}$ is the pmf of $\tilde{F}$, and we define $\Delta^{t}(\ell)\coloneqq\tilde{F}(\ell)^{t}-\tilde{F}(\ell-1)^{t}$ for any $t\in[n]$, with $\Delta^{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ under the convention $0^{0}=0$.
Additional Notations.
For any vector $\bm{x}=(x_{1},x_{2},\dots,x_{n})\in[K]^{n}$ and any $i,j\in[n]$, we define $\bm{x}_{i:j}\coloneqq(x_{i},x_{i+1}\dots,x_{j})$. Further define the maxima $\max\bm{x}_{i:j}=\max_{i\leq t\leq j}x_{t}$ and $\max\bm{x}=\max_{t\in[n]}x_{t}$. If $i>j$, we adopt the convention $\max\bm{x}_{i:j}=1$ consistent with the assumption that $x_{t}\geq 1$ for all $t$.
8.1 Family of Candidate Prior Distributions
Given the predicted prior $\tilde{F}$ over $[K]$, we consider a family of candidate prior distributions by conditioning on having values at most $k\in[K]$.
Definition 7.
Let $\mathcal{F}=\{F_{k}\}_{k\in[K]}$ be a family of distributions such that $F_{k}$ is the conditional distribution obtained by restricting $\tilde{F}$ to support $[k]$. In other words, the cdf of $F_{k}$ is
$F_{k}(\ell)~=~\begin{cases}\dfrac{\tilde{F}(\ell)}{\tilde{F}(k)},&1\leq\ell\leq k~;\\[8.61108pt] ~1,&k<\ell\leq K~.\end{cases}$
At the two extremes, we have $F_{K}=\tilde{F}$ and $F_{1}$ being a point mass at $1$. We use $\operatorname{\mathbf{Pr}}_{k}$ to denote probability under distribution $F_{k}$, and $\operatorname{\mathbf{Pr}}$ for probabilities of events independent of the specific choice of the underlying distribution, e.g., those over the randomness of the algorithm.
The main property of this construction is the following. Given any realized values in the first $t$ steps, the only information we can conclude about the real prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Formally, we have the next lemma.
Lemma 16.
For any $t\in[n]$, any $\ell\in[K]$, and any subset $S\subseteq[\ell]^{t}$ of $t$ realized values whose maximum equals $\ell$, the conditional probability:
$\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\mid\max\bm{x}_{1:t}=\ell\big]$
is the same for all $k$ with $\ell\leq k\leq K$.
Proof.
By Bayes’ rule, we have
$\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\mid\max\bm{x}_{1:t}=\ell\big]~=~\frac{\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\,,\,\max\bm{x}_{1:t}=\ell\big]}{\operatorname{\mathbf{Pr}}_{k}\big[\max\bm{x}_{1:t}=\ell\big]}~.$
Replacing $k$ with $K$ corresponds to multiplying both the numerator and denominator by $\tilde{F}^{t}(k)$ by Definition 7, resulting in the same ratio. ∎
8.2 Representations of Algorithms
To derive the linear constraints in polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$, we consider two representations of online algorithms when the predicted prior is $\tilde{F}$ and the real prior is one of the conditional distributions in the above family $\mathcal{F}$. The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., buchbinderSecretaryProblemsLinear2014, correaSampledrivenOptimalStopping2021).
Definition 8 (Conditional Acceptance Probabilities).
Given any algorithm, and for any $t\in[n]$, define $\textsc{Acc}_{t}:[K]^{t}\to[0,1]$ as the probability of accepting the value at time $t$, conditioned on the realized values up to time $t$, and not accepting any value before $t$. That is
$\textsc{Acc}_{t}(\bm{x}_{1:t})~\coloneqq~\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,\bm{x}_{1:t}=\bm{x}_{1:t}\,\big]~.$
The conditional acceptance probabilities fully characterize the algorithm, as we can write the probability of accepting the value at step $t$ as:
$\operatorname{\mathbf{Pr}}\big[\textsc{Alg}=t\mid\bm{x}\big]~=~\textsc{Acc}_{t}(\bm{x}_{1:t})\prod_{i\in[t-1]}\big(1-\textsc{Acc}_{i}(\bm{x}_{1:i})\big)~.$
Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a best-so-far value $x_{t}$ at time $t$ is independent of smaller values $\bm{x}_{1:t-1}$ in the past.
Definition 9 (Minor-oblivious Algorithms).
An algorithm is called minor-oblivious if $\textsc{Acc}_{t}(\bm{x}_{1:t})=\textsc{Acc}_{t}(\bm{y}_{1:t})$ for any $\bm{x}_{1:t}$ and $\bm{y}_{1:t}$ satisfying $x_{t}=y_{t}=\max\bm{x}_{1:t}=\max\bm{y}_{1:t}$. We then write $\textsc{Acc}_{t}(\ell)$ for the common value of $\textsc{Acc}_{t}(\bm{x}_{1:t})$ for all $\bm{x}_{1:t}$ satisfying $\max\bm{x}_{1:t}=\ell$.
Recall that based on the realization of $\bm{x}_{1:t}$, the only information we may conclude about the prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Intuitively, there is no reason to make decisions based on the details of $\bm{x}_{1:t}$ other than a maximum $\max\bm{x}_{1:t}$. In this sense, minor-oblivious algorithms are without loss of generality. We formalize this characterization in the following proposition, which may be of independent interest. The proof is deferred to Section C.1.
Proposition 17.
If an algorithm is $\alpha$-competitive when the prior is $F_{K}$ and $\beta$-competitive when the prior is $F_{k}$ for $k\in[K-1]$, there is a minor-oblivious algorithm with the same guarantees.
Focusing on minor-oblivious algorithms, we next consider a new representation by the probability of rejecting the first $t$ values conditioned on a maximum among them.
Definition 10 (Minor-oblivious Rejection Probabilities).
For any $t\in[n]$ and any $\ell\in[K]$, define the minor-oblivious rejection probability $\textsc{Rej}_{t}:[K]\to[0,1]$ as:
$\textsc{Rej}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ to simplify some expressions in the rest of the section.
There is a one-to-one correspondence between the two representations (see Section C.2). We state below the direction to be used in the proof of Theorem 15.
Lemma 18.
For any minor-oblivious rule, any $t\in[n]$, and any $\ell\in[K]$, we have
$\textsc{Acc}_{t}(\ell)~=~1-\frac{\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t}=\ell\,\big]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]}{\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t-1}=m,x_{t}=\ell\,\big]}~.$ (12)
8.3 Proof of Theorem 15
Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not best-so-far, since we consider MaxProb. We next verify that $y_{t,\ell}=\textsc{Rej}_{t}(\ell)$ is feasible w.r.t. polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$.
Bounds for Rejection Probabilities.
By the definition of the minor-oblivious rejection probabilities, and the artificial definition of $\textsc{Rej}_{0}(\ell)$, we obtain Constraints (6) and (7):
$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,\qquad\forall t\in[n],\;\forall\ell\in[K]~.$
Bounds for Acceptance Probabilities.
By $0\leq\textsc{Acc}_{t}(\ell)\leq 1$ and Equation 12, we get that
$0$ $~\leq~\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t}=\ell]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell,x_{t}<\ell]$
$~\leq~\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m,x_{t}=\ell]~.$
These two inequalities correspond to Constraints (8) and (9), because
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t}=\ell]$ $~=~\Delta^{t}(\ell)~,$
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell,x_{t}<\ell]$ $~=~\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell]\cdot\operatorname{\mathbf{Pr}}_{K}[x_{t}<\ell]$
$~=~\Delta^{t-1}(\ell)\tilde{F}(\ell-1)~,$
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m,x_{t}=\ell]$ $~=~\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m]\cdot\operatorname{\mathbf{Pr}}_{K}[x_{t}=\ell]$
[…]
Competitive Ratios.
Consider the probability of accepting a maximum value when the prior is $F_{k}$. We write it as:
$\operatorname{\mathbf{Pr}}_{k}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]$
$~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{k}\big[\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not best-so-far.
By the relation between $\tilde{F}=F_{K}$ and $F_{k}$, we can rewrite it as:
$\frac{1}{\tilde{F}(k)^{n}}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\max\bm{x}_{1:t}=\ell\,\big]$
$\qquad\qquad~=~\frac{1}{\tilde{F}(k)^{n}}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]~.$
Further, the event $\textsc{Alg}=t$ only depends on $\bm{x}_{1:t}$, while the realization of $\bm{x}_{t+1:n}$ is independent to both $\bm{x}_{1:t}$ and whether $\textsc{Alg}=t$. We can further transform the above expression into:
$\frac{1}{\tilde{F}^{n}(k)}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\ell\,\big]\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{t+1:n}\leq\ell\,\big]$
$\qquad\qquad~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~.$ (13)
The final step is to write the above probability as a linear expression of the minor-oblivious rejection probabilities $\textsc{Rej}_{t}(\ell)$’s. First, we write $\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$ as:
$\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t}=\ell\,\big]\,-\,\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~.$ (14)
The first part equals:
$\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}\leq\ell\,,\,x_{t}=\ell\,\big]+\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}=\ell\,,\,x_{t}<\ell\,\big]$
$\qquad~=~\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}[\max\bm{x}_{1:t-1}=m]\cdot\operatorname{\mathbf{Pr}}[x_{t}=\ell]\cdot\operatorname{\mathbf{Pr}}\big[\textsc{Alg}>t-1\mid\max\bm{x}_{1:t-1}=m\big]$
$\qquad\qquad\qquad~+\operatorname{\mathbf{Pr}}[\max\bm{x}_{1:t-1}=\ell]\cdot\operatorname{\mathbf{Pr}}[x_{t}<\ell]\cdot\operatorname{\mathbf{Pr}}\big[\textsc{Alg}>t-1\mid\max\bm{x}_{1:t-1}=\ell\big]$
$\qquad~=~\sum_{m\in[\ell]}\Delta^{t-1}(m)\cdot\tilde{f}(\ell)\cdot\textsc{Rej}_{t-1}(m)+\Delta^{t-1}(\ell)\cdot\tilde{F}(\ell-1)\cdot\textsc{Rej}_{t-1}(\ell)~.$ (15)
The second part equals:
$\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\ell\,\big]\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\mid\max\bm{x}_{1:t}=\ell\,\big]~=~\Delta^{t}(\ell)\cdot\textsc{Rej}_{t}(\ell)~.$ (16)
Combining Equations 13, 14, 15 and 16, we conclude that when the real prior is $F_{k}$, the competitive ratio is equal to
$\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\bigg($ $\sum_{m\in[\ell]}\Delta^{t-1}(m)\cdot\tilde{f}(\ell)\cdot\textsc{Rej}_{t-1}(m)$
$\quad+\Delta^{t-1}(\ell)\cdot\tilde{F}(\ell-1)\cdot\textsc{Rej}_{t-1}(\ell)-\Delta^{t}(\ell)\cdot\textsc{Rej}_{t}(\ell)\bigg).$
Since the algorithm is $\alpha$-consistent and $\beta$-robust, the above competitive ratio is at least $\alpha$ when $k=K$, and at least $\beta$ when $k\in[K-1]$. These correspond to Constraints (10) and (11).
8.4 Numerical Results
For any $\lambda\in[0,1]$, we solve a linear program with variables $\{y_{t,\ell}\}_{t\in[n],\,\ell\in[K]}\cup\{\alpha,\beta\}$ to maximize:
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta$
subject to the linear constraints that define polytope $\mathcal{P}_{n,K,\,\tilde{F}}(\alpha,\beta)$. The optimal objective value $\mathrm{LP}^{*}(\lambda)$ then implies a hardness result of the form:
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta~\geq~\mathrm{LP}^{*}(\lambda)~.$
We set $n=30$ and $K=1024$, and consider a predicted prior $\tilde{F}$ defined by $\tilde{f}(k)\propto\frac{1}{k}$ for $k\in[K]$. We solve the linear programs for $\lambda=\frac{i}{100}$ for $0\leq i\leq 100$ to get the curve shown in Figure 1(b). For details, see our code repository.
Appendix A Missing Proofs for MaxProb Consistency
A.1 Monotonicity of Consistency
Lemma 19.
For any fixed $\lambda_{1},\lambda_{2}\in[0,1]$, the consistency $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is non-increasing in $n$.
Proof.
We show that for any $n$, it holds:
$\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\leq~\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~.$
Consider a relaxed version of the problem with $n+1$ values, where the adversary tells the algorithm at the beginning the information about the smallest realized value, including its arrival time $\tau$ and its value $x_{\tau}$. We consider two policies.
In the first policy, we ignore the additional information and directly apply the algorithm parameterized by $\theta^{*}_{n+1}$. By definition, the consistency of this policy is $\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
We now describe the second and optimal policy. Conditioned on $\tau$ and $x_{\tau}$, the remaining $n$ values are i.i.d. from $F$ restricted to $x_{\tau},+\infty)$, and their arrival times are i.i.d. uniform on $[0,1]$. Thus, by the optimality of $\theta^{*}_{n}$, any algorithm that rejects all values during $0,\lambda_{1}]$, accepts best-so-far values during $\lambda_{1},\lambda_{2}]$ based on a threshold function, and accepts best-so-far values during $\lambda_{2},1]$ has consistency at most $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
Since the second strategy is optimal, we get the stated inequality. ∎
A.2 Asymptotic Characterization of $\theta_{n}^{*}$
Lemma 20.
For any fixed $s\in[0,1)$, we have
$\theta^{*}_{n}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
where $\gamma\approx 0.80435$ satisfies
$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$
Proof.
Recall Equation 5:
$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta^{*}_{n}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$
Fix any $s$, substitute $x_{n}=\frac{1}{\theta^{*}_{n}(s)}-1$, the left-hand side becomes
$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
Expanding the left-hand side with the binomial theorem, we have
$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k!}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{(n-1)!}x_{n}^{n-1}=1~.$
Now, we analyze the asymptotic behavior of $x_{n}$. Obviously, we have $x_{n}=\mathcal{O}(\tfrac{1}{n-1})$; otherwise, the left-hand side either diverges to infinity or converges to 0, thus cannot equal $1$.
Now, we find the constant such that $(n-1)(1-s)x_{n}\to c$ for some constant $c$ as $n\to\infty$. Note that such a constant $c$ shall satisfy
$c+\frac{n-2}{n-1}\cdot\frac{c^{2}}{2!2}+\cdots+\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\cdot\frac{c^{k}}{k!k}+\cdots=1.$
Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that $c$ satisfies
$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
which defines the constant $\gamma\approx 0.80435$ in the lemma.
We can prove the convergence more rigorously by showing that, for any $c\in(0,1)$, the difference between the ideal and the actual sum tends to 0 as $n\to\infty$.
Consider the difference:
$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\left(1-\left(1-\frac{1}{n-1}\right)\left(1-\frac{2}{n-1}\right)\cdots\left(1-\frac{k-1}{n-1}\right)\right)~.$
Observe that $1\leq(1-x_{1})(1-x_{2})\cdots(1-x_{m})+x_{1}+x_{2}+\cdots+x_{m}$ for any $x_{i}\in[0,1]$. Indeed, the partial derivative shows that the right-hand side is increasing in each $x_{i}$, thus the minimum is attained at $x_{i}=0$ for all $i$, which is $1$.
It follows that the difference is bounded by
$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{1+2+\cdots+(k-1)}{n-1}=\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{k(k-1)}{2(n-1)}\leq\frac{1}{2(n-1)}\sum_{k=1}^{\infty}\frac{1}{(k-1)!}~.$
Using the fact that $\sum_{k=1}^{\infty}\frac{1}{(k-1)!}=\mathrm{e}$, the bound is $\mathcal{O}(\tfrac{1}{n})$ and therefore tends to 0 as $n\to\infty$.
Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$. This completes the proof by $\theta^{*}_{n}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$. ∎
Appendix B Numerical Solution to the Differential Equation in Theorem 1
Recall the differential equation in Theorem 1:
$\theta(\lambda_{1})$ $~=~1~,$ (17)
$\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\theta\big(\max\{s,z\}\big)^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall z\in[\lambda_{1},\lambda_{2}]~.$ (18)
We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in[\lambda_{2},1]$.
Define the left-hand side of Equation 18 as $L(z)$ and simplify:
$L(z)~:=$ $~\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\theta\big(\max\{s,z\}\big)^{t}\,\mathrm{d}s\,\mathrm{d}t$ (19)
$~=$ $~\int_{z}^{1}\bigg(\frac{z\,\theta(z)^{t}}{t}+\int_{z}^{t}\frac{\theta(s)^{t}}{t}\,\mathrm{d}s\bigg)\,\mathrm{d}t$
$~=$ $~z\int_{z}^{1}\frac{\theta(z)^{t}}{t}\,\mathrm{d}t+\int_{z}^{\lambda_{2}}\int_{s}^{1}\frac{\theta(s)^{t}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $\lambda_{2},1]$.
We approximate $\theta$ by a left‑continuous non‑increasing step function computed backward from $\lambda_{2}$ to $\lambda_{1}$ so that Equation 18 holds for all $z\in[\lambda_{1},\lambda_{2}]$.
Definition 11 (Step function approximation).
Given an integer $m>0$, define a uniform grid $\{z_{i}\}_{i=1}^{m+1}$ by $z_{i}=\lambda_{1}+\tfrac{i-1}{m}(\lambda_{2}-\lambda_{1})$ for $i=1,\dots,m+1$ (so $z_{m+1}=\lambda_{2}$). We say that $\theta$ is a step function with step values $\theta_{1},\dots,\theta_{m}$ if $\theta(z)=\theta_{i}$ for all $z\in(z_{i},z_{i+1}]$, and $\theta(\lambda_{1})=\theta_{1}$.
Lemma 21 (Monotonicity of $L(z)$).
$L(z)$ is decreasing in $z$ on $\lambda_{1},\lambda_{2}]$.
Proof.
As $z$ increases, the integration domain in Equation 19 shrinks with a non-negative integrand. The integrand is non-increasing since $\theta$ is non-increasing. ∎
Lemma 22 (Endpoint reduction).
A step function $\theta$ with $m$ steps satisfies Equation 18 for all $z\in[\lambda_{1},\lambda_{2}]$ if it satisfies Equation 18 at the endpoints of each interval, i.e., for all $i\in[m]$,
$L(z_{i+1})~\geq~\alpha\,\theta_{i}~.$
Proof.
By Lemma 21, $L(z)$ is decreasing in $z$. Thus, for any $z\in(z_{i},z_{i+1}]$, it holds $L(z)\geq L(z_{i+1})\geq\alpha\,\theta_{i}=\alpha\,\theta(z)$. ∎
Next, we describe how to compute the step values $\theta_{1},\dots,\theta_{m}$ given $\alpha$ and $m$ so that Lemma 22 holds.
Initialization at $z=\lambda_{2}$.
For $i=m$, Lemma 22 is
$\lambda_{2}\int_{\lambda_{2}}^{1}\frac{\theta(\lambda_{2})^{t}}{t}\,\,\mathrm{d}t~\geq~\alpha\,\theta(\lambda_{2})~.$
Letting $\theta_{m}\coloneqq\theta(\lambda_{2})\neq 0$, this is
$\lambda_{2}\int_{\lambda_{2}}^{1}\frac{\theta_{m}^{-(1-t)}}{t}\,\,\mathrm{d}t~\geq~\alpha~,$
where the left-hand side decreases from $\infty$ to 0 as $\theta_{m}$ increases from 0 to $\infty$. Thus, there exists a $\theta_{m}^{*}>0$ such that the equality holds. [^9] During experiment, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
Backward recursion.
For $i=m-1,m-2,\ldots,1$, suppose we have computed $\theta_{i+1}^{*},\theta_{i+2}^{*},\ldots,\theta_{m}^{*}$. Using Equation 19, we have
$L(z_{i+1})~=$ $z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\int_{z_{j}}^{z_{j+1}}\int_{s}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\,\,\mathrm{d}s$
$~=$ $z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta_{j}^{*}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~.$
Thus, Lemma 22 is simplified to
$z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta_{j}^{*}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~\geq~\alpha\,\theta_{i}~.$ (20)
Equivalently,
$z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{-(1-t)}}{t}\,\,\mathrm{d}t+\theta_{i}^{-1}\,\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta^{*}_{j}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~\geq~\alpha~,$ (21)
whose left-hand side decreases from $\infty$ to 0 as $\theta_{i}$ increases from 0 to $\infty$. Again, there exists a $\theta_{i}^{*}>0$ such that the equality holds.
We select the largest feasible $\theta_{i}=\theta_{i}^{*}$ by binary search such that equality holds.
Lemma 23 (Monotonicity of the step values).
The sequence $(\theta_{i}^{*})_{i=1}^{m}$ produced by the recursion is non-increasing in $i$.
Proof.
By definition, $\alpha\theta_{i}^{*}=L(z_{i+1})\geq L(z_{i+2})=\alpha\theta_{i+1}^{*}$, where the inequality holds by Lemma 21. ∎
Enforcing $\theta(\lambda_{1})=1$.
We only need to find an $\alpha$ with the corresponding sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfying Equation 18 and $\theta_{1}^{*}\geq 1$. The following lemma shows this suffices.
Lemma 24.
If a step function $\theta$ with step values $\theta_{1},\theta_{2},\dots,\theta_{m}$ satisfies Equation 18, then the step function with step values $(\min\{\theta_{i},1\})_{i=1}^{m}$ also satisfies Equation 18.
Proof.
By unifying intervals that $\theta_{i}$ exceeds $1$, as Lemma 22, we only need to check that Equation 21 holds for $\min\{\theta_{i},1\}$, where $i$ is the last $\theta_{i}$ exceeding $1$. This instantly stands by the non-increasing of the left-hand side. ∎
A direct corollary implies the desired result.
Corollary 25.
If an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 17 and Equation 18.
In our implementation, we fix several pairs of values $(\alpha,\beta)$ and numerically verify that for each pair, the constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$. Then Corollary 25 implies that the algorithm $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ is $\alpha$-consistent and $\beta$-robust. The results are plotted in Figure 1(a).
Intuitively, as $m\to\infty$, the step function approximation tightens. In experiments, we set $m=300$. For details, see the implementation of the backward recursion as well as the solved $\theta_{i}^{*}$’s in our code repository.
Figure 3 illustrates the case when $\beta=0.01$ and $m=300$ yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$. This implies that $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ with $\theta^{*}$ being the step function with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{300}$ is an $\alpha\approx 0.6908$-consistent and $\beta=0.01$-robust algorithm.
Figure 3: Example solution of $\theta$ when $\beta=0.01$, $m=300$, and $\alpha\approx 0.6908$.
Appendix C Missing Proofs for MaxProb Hardness
C.1 Minor-Oblivious Algorithms: Proof of Proposition 17
The key observation is that, when the current value is best-so-far and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability. This is formalized in the following lemma.
Lemma 26.
For any $t\in[n]$ and any $\ell\in[K]$, the conditional probabilities
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t,\ x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$
are identical for all $k$ with $\ell\leq k\leq K$.
Proof.
By Bayes’ rule, we can obtain that
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\frac{\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\,,\,x_{t}=\ell\mid\max\bm{x}_{1:t}=\ell\,\big]}{\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\ell\mid\max\bm{x}_{1:t}=\ell\,\big]}~.$
Denote by $S\subseteq[K]^{t}$ the set of all $\bm{x}_{1:t}$ with $x_{t}=\max\bm{x}_{1:t}=\ell$. The numerator can be written as
$\sum_{\bm{x}_{1:t}\in S}\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\bm{x}_{1:t}\,\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[\,\bm{x}_{1:t}\mid\max\bm{x}_{1:t}=\ell\,\big]~,$
where the second factor is the conditional probability of realizing $\bm{x}_{1:t}$ given $\max\bm{x}_{1:t}=\ell$. This is invariant for all $k$ with $\ell\leq k\leq K$ by Lemma 16. The same holds for the denominator. Hence, the ratio is invariant. ∎
Given any algorithm $\mathcal{A}$, whose index of accepted value is Alg, we construct a minor-oblivious algorithm $\mathcal{A}^{\prime}$ with conditional acceptance probabilities
$r^{\prime}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t,\ x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~.$
That is, the probability of accepting a best-so-far value $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is best-so-far; and 3) $\mathcal{A}$ rejects the first t1t-1 values. By Lemma 26, we may define this conditional probability with respect to prior $F_{K}$ without loss of generality.
Finally, we show that algorithm $\mathcal{A}^{\prime}$ accepts the maximum value with the same probability as $\mathcal{A}$. Let $\textsc{Alg}^{\prime}$ be the index of the accepted value by algorithm $\mathcal{A}^{\prime}$. For any prior $F_{k}$, observe that
$\operatorname{\mathbf{Pr}}_{k}\left[x_{\textsc{Alg}}=x_{\textsc{Opt}}\right]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\,\operatorname{\mathbf{Pr}}_{k}\left[\max\bm{x}_{t+1:n}\leq\ell\right]\cdot\operatorname{\mathbf{Pr}}\nolimits_{k}\left[\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]$
$\qquad\qquad\qquad\cdot\operatorname{\mathbf{Pr}}\left[\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]~.$
Similarly, we have
$\operatorname{\mathbf{Pr}}_{k}\left[x_{\textsc{Alg}^{\prime}}=x_{\textsc{Opt}}\right]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\,\operatorname{\mathbf{Pr}}_{k}\left[\max\bm{x}_{t+1:n}\leq\ell\right]\cdot\operatorname{\mathbf{Pr}}\nolimits_{k}\left[\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]$
$\qquad\qquad\qquad\cdot\operatorname{\mathbf{Pr}}\left[\textsc{Alg}^{\prime}=t\mid\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]~.$
Comparing the two expressions, the first parts coincide, and the third parts are equal by the construction of $\mathcal{A}^{\prime}$. It remains to prove that for any $t\in[n]$ and any $k,\ell\in[K]$:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~.$ (22)
We next prove this by a joint induction with the following claim:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}^{\prime}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]~=~\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]~.$ (23)
The base case of $t=1$ holds trivially, as the probabilities are independent of the algorithms’ decisions. Next, suppose that both equalities hold for some $t\geq 1$. We will prove them for $t+1$. Enumerating over the state at step $t$, we have:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t+1\,,\,x_{t+1}=\max\bm{x}_{1:t+1}=\ell\,\big]$
$\quad~=~\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=m\,\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[x_{t+1}=\ell\big]$ (24)
$\qquad\qquad\cdot\,\big(1-\operatorname{\mathbf{Pr}}[\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=m]\big)$ (25)
$\quad\qquad+\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}\nolimits_{k}\big[\textsc{Alg}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=m\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[x_{t+1}=\ell\big]\,.$ (26)
By the induction hypothesis, we can replace Alg with $\textsc{Alg}^{\prime}$ in (24) and (26) while preserving the value. Further, we can also replace Alg with $\textsc{Alg}^{\prime}$ in (25) by the construction of algorithm $\mathcal{A}^{\prime}$. Hence, Equation 22 holds for $t+1$ as well.
The proof of Equation 23 for $t+1$ is almost verbatim.
C.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 18
From Rejection Probabilities to Acceptance Probabilities.
By definition, we have:
$\textsc{Acc}_{t}(\ell)~=$~ $\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$
$~=$~ $1-\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]\,.$
By Bayes’ rule, the probability on the right-hand side can be written as
$\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\frac{\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]}{\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\mathbf{x}_{1:t}=\ell\,\big]}\,.$
Here we consider the prior $F_{K}$ without loss of generality, because using any other $F_{k}$ multiplies numerator and denominator by the same $\tilde{F}(k)^{-t}$ factor according to Definition 7.
The numerator further equals
$\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~-~\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]$
$\quad~=~\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\ell\,\big]\cdot\textsc{Rej}_{t}(\ell)-\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]\cdot\textsc{Rej}_{t-1}(\ell)~.$
The denominator equals
$\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}=m\,,\,x_{t}=\ell\,\big]=\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t-1}=m\,,\,x_{t}=\ell\,\big]\cdot\textsc{Rej}_{t-1}(m)\,.$
Finally, putting these together proves Lemma 18.
From Acceptance Probabilities to Rejection Probabilities.
We remark that the relation between acceptance probabilities and rejection probabilities is one-to-one. Given the acceptance probabilities, we can inductively deduce the rejection probabilities using Lemma 18. Restating Equation 12 from Lemma 18:
$\textsc{Acc}_{t}(\ell)~=~1-\frac{\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t}=\ell\,\big]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]}{\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t-1}=m,x_{t}=\ell\,\big]}\,.$
Given $\textsc{Rej}_{t-1}(\cdot)$ and $\textsc{Acc}_{t}(\cdot)$ for any $t\geq 1$, we can derive $\textsc{Rej}_{t}(\cdot)$ from the equation above.

### v2

@@
-Optimal Stopping with a Predicted Prior
-Optimal Stopping with a Predicted Prior
+
+Optimal Stopping with a Predicted Prior [^1] This second arXiv version improves the exposition based on feedback we received after posting the first version.
+We acknowledge the concurrent work by Kehne and Kesselheim [40], add Section 3.3 that explains how to use the indifference Bellman equation to compute the optimal threshold function within our algorithm class, and strengthen Theorem 1 by establishing that the trade‑off curve for MaxProb is tight within the class.
 Tian Bai [^2] University of Bergen.
@@
 Dongchen Li22footnotemark: 2
-(November 2025)
 There are two major models of value uncertainty in the optimal stopping literature: the secretary model, which assumes no prior knowledge, and the prophet inequality model, which assumes full information about value distributions.
@@
 The decision maker can accept at most one value and aims to maximize its expectation.
-An optimal solution is Dynkin’s algorithm [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989], which rejects the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values, and then accepts the first value that exceeds all previous ones.
-It guarantees that the expected accepted value is at least $\frac{1}{\mathrm{e}}$ times the maximum.
+An optimal solution is Dynkin’s algorithm [21, 29], which rejects the first $\lfloor\nicefrac{{n}}{{\mathrm{e}}}\rfloor$ values, and then accepts the first value that exceeds all previous ones.
+It guarantees that the expected accepted value is at least $\nicefrac{{1}}{{\mathrm{e}}}$ times the maximum.
 This ratio of the algorithm’s expected accepted value to the maximum value in hindsight is called the competitive ratio.
-In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from known prior distributions.
+In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from fully known prior distributions.
 In the i.i.d.
-case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [hillComparisonsStopRule1982, kertzStopRuleSupremum1986, correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021], substantially better than prior-free secretary guarantee of $\frac{1}{\mathrm{e}}\approx 0.368$.
+case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [38, 41, 16, 44], substantially better than the prior-free secretary guarantee of $\nicefrac{{1}}{{\mathrm{e}}}\approx 0.368$.
 On the one hand, the secretary problem does not capture the abundant data available to decision makers in this era.
@@
 What real-world decision makers often have are machine-learned and potentially erroneous predicted priors.
-duttingPostedPricingProphet2019 explored this direction for the general prophet inequality problem.
-They examined how mispredicted priors affect existing methods and showed a degradation in performance proportional to the prediction error.
-This can be viewed as an analysis of the smoothness of algorithms under the framework of algorithms with predictions [mitzenmacher2022algorithms].
+Dütting and Kesselheim [20] took a step in this direction.
+They show that the performance of the optimal backward-induction algorithm that fully trusts the predicted prior degrades proportionally to some appropriate measures of prediction error, thereby providing a smoothness-style guarantee in the framework of algorithms with predictions [46].
 1.1 Optimal Stopping with a Predicted Prior
@@
 Such an algorithm leverages prior knowledge when the predicted prior is correct, while retaining meaningful worst-case guarantees.
-The $\frac{1}{\mathrm{e}}$-competitive Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust for both objectives.
-The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by gilbertRecognizingMaximumSequence are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent respectively, but only 0-robust (see Section 2).
+The $\nicefrac{{1}}{{\mathrm{e}}}$-competitive Dynkin’s algorithm is $\nicefrac{{1}}{{\mathrm{e}}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust for both objectives.
+The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [38, 41] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by Gilbert and Mosteller [33] are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent, respectively, but only 0-robust (see Section 2).
 Further, randomly running either Dynkin’s algorithm or an i.i.d.
 prophet inequality algorithm linearly interpolates between their consistency and robustness.
-Finally, the optimality of these competitive ratios implies that no algorithm is better than $\frac{1}{\mathrm{e}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb.
+Finally, the optimality of these competitive ratios implies that no algorithm is better than $\nicefrac{{1}}{{\mathrm{e}}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb.
 These baseline algorithms and impossibilities leave a significant gap between them.
@@
 Bi-criteria Algorithms.
-We define a family of bi-criteria algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$.
+We define a family of bi-criteria algorithms $\mathcal{A}(\tilde{F},\theta)$ parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$.
 Upon observing each value and when no value has been accepted, a bi-criteria algorithm accepts the current value if:
   1. 1.
-    It is best-so-far, i.e., greater than or equal to all previous values; (prior-free criterion)
+    It is a record, i.e., greater than or equal to all previous values; (prior-free criterion)
   2. 2.
     Its cdf value w.r.t. $\tilde{F}$ exceeds a time-dependent threshold $\theta(t)$. (prior-dependent criterion)
-For now, readers may interpret $t$ as the time step by which the value is observed. Later, we will also consider a continuous-time model to simplify the analysis (see Section 2 for a discussion).
+For now, readers may interpret $t$ as the time step by which the value is observed.
+Later, we will also consider an equivalent continuous-time model to simplify the analysis (see Section 2 for a discussion).
 We set the thresholds as follows.
 Before some time $\lambda_{1}$, the threshold is $1$, and the algorithm rejects all values.
-After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free best-so-far criterion.
+After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free record-based criterion.
 These two prior-free phases yield robustness.
@@
 Theorem 1 (MaxExp algorithm).
-Suppose $0\leq\beta\leq\frac{1}{e}$ and $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
-Further, suppose there is a left-continuous non-increasing function $\theta:[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
-$\theta(\lambda_{1})$ $~=~1~,$
-$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall\lambda_{1}\leq z\leq\lambda_{2}~.$
-Then, there is an $\alpha$-consistent $\beta$-robust algorithm for the MaxExp objective.
-To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation in Theorem 1, which does not appear to admit a closed-form solution. See Appendix B.
+For any $\beta\in[0,\nicefrac{{1}}{{e}}]$, suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
+If there is a left-continuous non-increasing function $\theta\colon[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
+$\theta(\lambda_{1})$ $~=~1~,$ (1)
+$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall z\in[\lambda_{1},\lambda_{2}]~,$
+then for any $n\in\mathbb{N}$, there is a bi-criteria algorithm that is $\alpha$-consistent and $\beta$-robust for the MaxExp objective.
+The proof of Theorem 1 is given in Section 6.
+To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation (1), which does not appear to admit a closed-form solution.
+See Appendix B.
 Theorem 2 (MaxProb algorithm).
-Suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
-There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective for any $\beta\in[0,\frac{1}{e}]$ and
+For any $\beta\in[0,\frac{1}{e}]$, suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
+For any $n\in\mathbb{N}$, there is an $\alpha$-consistent $\beta$-robust bi-criteria algorithm for the MaxProb objective, with
 $\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\,\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
 where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
-We further prove that certain consistency-robustness trade-offs cannot be achieved for the MaxProb objective. In particular, there is no “best-of-both-world” algorithm.
+Moreover, this trade-off is tight within the class of bi-criteria algorithms in the sense that, for any specified $(\alpha,\beta)$, none of the bi-criteria algorithms can improve both objectives simultaneously.
+The proof of Theorem 2 is given in Section 7.
+For the MaxProb objective, besides the tightness within the bi-criteria family above, we further show a non-trivial hardness that certain consistency-robustness trade-offs cannot be achieved by any algorithm.
+In particular, this implies that there is no “best-of-both-world” algorithm for this objective.
 Theorem 3 (MaxProb hardness, informal version of Theorem 17).
 There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective only if a polytope $\mathcal{P}(\alpha,\beta)$ is feasible.
-In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
+In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust.
+Again, to plot the curve in Figure 1(b), we rely on numerical linear program solvers to solve the feasibility of certain polytopes, which does not appear to admit a closed-form solution.
 1.3 Overview of Techniques
@@
 By definition, we may consider a real prior $F$ and a (mis)predicted prior $\tilde{F}$ that misleads the algorithm.
-Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq(F\circ\tilde{F}^{-1}\circ\theta)(t)$.
+Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq F\circ\tilde{F}^{-1}\circ\theta(t)$.
 The latter viewpoint is much easier to work with — we just need to restrict the misspecified threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ via our choice of threshold $\theta$.
-Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ has the same value of 0 or $1$.
-This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ is between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any value that is best-so-far.
+Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ should have the same value of 0 or $1$.
+This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ lies between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any record.
 See Section 3 for a detailed discussion, and Section 5 for the robustness analysis.
 Implicit Sharding.
-We show that optimal stopping with a predicted prior is hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest.
-This is inspired by the sharding technique by harbNewProphetInequalities2025, which breaks each value into many i.i.d.
+We show that optimal stopping with a predicted prior is the hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest.
+This is inspired by the sharding technique by Harb [37], which breaks each value into many i.i.d.
 shards based on the prior, such that the original value equals the maximum shard.
@@
 Prophet Inequalities with Full Prior Information.
-hillComparisonsStopRule1982 proposed a $(1-\frac{1}{\mathrm{e}})$-competitive algorithm for i.i.d.
-prophet inequality and the MaxExp objective.
-This was improved to $0.738$ by abolhassaniBeating11Ordered2017, and finally to approximately $0.745$ by correaPostedPriceMechanisms2017.
-The latter ratio is tight, matching a hardness result by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986].
-For general prophet inequality with non-identical distributions, the optimal competitive ratio is $0.5$ [krengel1978semiamarts, samuel1984comparison].
-The prophet secretary problem [esfandiariProphetSecretary2015] interpolates between general and i.i.d.
-prophet inequalities, with values drawn from non-identical distributions and arriving by a random order.
-The optimal competitive ratio for prophet secretary is open, with the best lower and upper bounds being $0.688$ [chenProphetSecretaryMatching2025] and $0.723$ [giambartolomeiProphetInequalitiesSeparating2023].
-For the MaxProb objective, gilbertRecognizingMaximumSequence gave an optimal algorithm for the i.i.d.
-case, while samuels1982 derived an explicit expression of the optimal competitive ratio of approximately $0.580$.
-For non-identical distributions, esfandiariProphetsSecretariesMaximizing2020 obtained an optimal $\frac{1}{\mathrm{e}}$ competitive ratio in the prophet inequality model, and recovered the $0.580$ ratio in the prophet secretary model under a “no-superstar” assumption.
-nutiSecretaryProblemDistributions2022 later removed this assumption.
-See correaRecentDevelopmentsProphet2019 for a survey. We stress that none of these prophet inequality algorithms provides non-trivial robustness.
+When the algorithm has complete knowledge of the distributions, the prophet inequality problem has been extensively studied.
+Under the MaxExp objective, Hill and Kertz [38] proposed a $(1-\nicefrac{{1}}{{\mathrm{e}}})$-competitive algorithm for i.i.d.
+prophet inequality.
+This was improved to $0.738$ by Abolhassani et al.
+[1], and finally to approximately $0.745$ by Correa et al.
+[16].
+The latter ratio is tight, matching a hardness result by Hill and Kertz [38, 41].
+For prophet inequality where values are independently drawn from non-identical distributions, the optimal competitive ratio is $0.5$ [43, 51].
+The prophet secretary problem [23] interpolates between these two settings: values are independently drawn from non-identical distributions and arrive in random order.
+The optimal competitive ratio for prophet secretary is open: a lower bound of $0.688$ [10] and an upper bound of $0.723$ [32] are currently known.
+For the MaxProb objective, Gilbert and Mosteller [33] gave an optimal algorithm for the i.i.d.
+prophet inequality, while Samuels [52] derived an explicit expression for the optimal competitive ratio of approximately $0.580$.
+For non-identical distributions, Esfandiari et al.
+[24] obtained an optimal $\nicefrac{{1}}{{\mathrm{e}}}$-competitive ratio for the prophet inequality model, and recovered the $0.580$ ratio for the prophet secretary under a “no-superstar” assumption.
+Nuti [47] later removed this assumption.
+See Correa et al. [15] for a survey. We stress that none of these prophet inequality algorithms provide non-trivial robustness guarantees.
 Prophet Inequalities with Partial Prior Information.
-There has been extensive research on the sample-based model, in which the algorithm has access to i.i.d.
-samples from the prior.
-One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [azar2014prophet, rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022, ezraProphetInequalitySamples2024].
-Notably, the optimal $\frac{1}{2}$-competitive prophet inequality can be achieved with only one sample from each distribution [rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022].
-Another line of work studies the sample complexity of learning a nearly optimal algorithm [diakonikolasLearningOnlineAlgorithms2021, correaUnknownIIDProphets2021, correaTwosidedGameGoogol2020, rubinsteinOptimalSinglechoiceProphet2020, cristiProphetInequalitiesRequire2023, guo2021generalizing, correaSampledrivenOptimalStopping2021, jin2024sample], culminating in the state-of-the-art $\tilde{\mathcal{O}}(\frac{n}{\varepsilon})$ sample complexity for i.i.d.
-prophet inequality [correaSampledrivenOptimalStopping2021], and $\tilde{\mathcal{O}}(\frac{1}{\varepsilon^{2}})$ sample complexity per distribution for the general case [jin2024sample].
-cristiProphetInequalitiesRequire2023 gave a unified proof that $\mathcal{O}(\mathrm{poly}(\frac{1}{\epsilon}))$ samples per distribution suffice for several major variants of prophet inequalities.
-Other models of partial information include learning nearly optimal algorithms via quantile queries [liProphetInequalityIID2023], bandit feedback [gatmiry2024bandit], or simple posted-pricing policies [perezsalazarIIDProphetInequality2023].
-These models and the sample-based model are incomparable to ours.
-They assume partial information from the correct prior, while the model of this paper considers a predicted prior that may be arbitrarily wrong.
+Relaxing the assumption of complete prior knowledge, several models have studied algorithms with partial information about the distributions.
+The most prominent one is the sample-based model, in which the algorithm has access to i.i.d.
+samples from the prior distribution(s).
+One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [4, 50, 35, 26].
+Notably, the optimal $\nicefrac{{1}}{{2}}$-competitive prophet inequality can be achieved with only one sample from each distribution [50, 35].
+Another line of work studies the minimal number of samples needed to achieve the optimal competitive ratio up to $\varepsilon$ [19, 13, 11, 50, 18, 36, 12, 39].
+The state-of-the-art bounds for i.i.d.
+prophet inequality are $\tilde{\mathcal{O}}(\nicefrac{{n}}{{\varepsilon}})$ for learning a $0.745-\varepsilon$ competitive algorithm [12], and $\tilde{\mathcal{O}}(\nicefrac{{n}}{{\varepsilon^{2}}})$ for learning an online policy that is optimal up to an $\varepsilon$ additive error [36].
+For general prophet inequality, $\tilde{\mathcal{O}}(\nicefrac{{1}}{{\varepsilon^{2}}})$ samples per distribution suffice [39].
+Finally, Cristi and Ziliotto [18] gave a unified proof that $\mathcal{O}(\mathrm{poly}(\nicefrac{{1}}{{\varepsilon}}))$ samples per distribution suffice for two variants of prophet inequalities.
+Other models of partial information include quantile-query oracles [28, 48] and bandit feedback [31].
+These models and the sample-based model are incomparable to ours: they assume access to the correct prior, while we consider a predicted prior that may be arbitrarily wrong.
 Algorithms with Predictions.
-Our work falls within the algorithms with predictions framework [mitzenmacher2022algorithms] for augmenting worst-case analysis of algorithms with machine-learned advice.
+Our work falls within the algorithms with predictions framework [46] for augmenting worst-case analysis of algorithms with machine-learned advice.
 Consistency, robustness, and smoothness are the primary performance metrics, capturing the algorithms’ performance when the prediction is correct, completely wrong, and partially wrong.
-The framework has been applied to many optimization problems, including scheduling [kumar2018improving] and caching [lykouris2021competitive].
-In the context of optimal stopping, antoniadisSecretaryOnlineMatching2020 studied a prediction of the maximum value, fujiiSecretaryProblemPredictions2024 assumed having a predicted sequence of values, and braun2024secretary considered a prediction of the gap between the maximum value and the $k$-th largest.
-In contrast, this paper considers distributional predictions, which aligns with the literature of prophet inequality and the machine-learned priors in practice.
-As mentioned earlier, duttingPostedPricingProphet2019 also considered mispredicted prior with a focus on smoothness, while this paper examines the trade-off between consistency and robustness.
+This framework has been applied to many online optimization problems, including ski-rental [49, 19, 6, 2], scheduling [49, 6], caching [45], online bidding [2], and so on.
+Within optimal stopping, prior work has explored several forms of predictions.
+Antoniadis et al.
+[3] studies a prediction of the maximum value, Fujii and Yoshida [30] and Balkanski et al.
+[5] assume having a predicted sequence of values, and Braun and Sarkar [7] considers a prediction of the gap between the maximum and the $k$-th largest value.
+In contrast, our work considers distributional predictions, bridging the extensive literature of prophet inequality with the growing use of machine-learned priors in practice.
+Within the context of prophet inequalities, the only relevant work to our knowledge is by Dütting and Kesselheim [20], which examines the impact of mispredicted priors with an emphasis on smoothness.
+This work is the first to investigate the consistency-robustness trade-off given a distributional prediction.
+In general, while distributional predictions have been explored in certain domains such as online matching, where they have been incorporated into traditional stochastic models [25, 9], this type of prediction remains largely underexplored.
+Independent and Concurrent Work.
+Kehne and Kesselheim [40] study the same problem, focusing on the MaxExp setting.
+Their results and ours are complementary.
+The algorithms in both papers are under the bi-criteria framework.
+They show a non-trivial consistency-robustness trade-off using a simple single-threshold between $\lambda_{1}$ and $\lambda_{2}$, whereas we optimize the choice of thresholds within the bi-criteria framework.
+They rule out “best-of-both-world” algorithms for the MaxExp setting, whereas we prove such an impossibility result for the MaxProb setting.
 2 Preliminaries
@@
 Let $\mathcal{D}$ be the set of all probability distributions supported on $0,\infty)$.
-We adopt the standard abbreviations cdf, pdf, and pmf for cumulative distribution function, probability density function, and probability mass function.
-We often use the same symbol for a distribution and its cdf (e.g., $F$), when clear from context.
+We adopt the standard abbreviations cdf, pdf, and pmf for the cumulative distribution function, probability density function, and probability mass function.
+We use the same symbol for a distribution and its cdf (e.g., $F$).
 Unless stated otherwise, expectations and probabilities are taken over the randomness of the values and any internal randomness of the algorithm.
@@
 Let $x_{\textsc{Opt}}$ denote the maximum value among all $n$ values, i.e., $x_{\textsc{Opt}}\coloneqq\max_{i\in[n]}x_{i}$.
-The max expectation (MaxExp) objective considers the expected value of the algorithm’s accepted decision, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$.
+The max expectation (MaxExp) objective considers the expectation of the algorithm’s accepted value, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$.
 The corresponding competitive ratio is
@@
 $\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~.$
-This coincides with the competitive ratio since the offline optimal choice is always maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
+This coincides with the competitive ratio since the offline optimal choice is always the maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
 2.2 Optimal Stopping with a Predicted Prior
@@
 Depending on what the algorithm knows about the prior distribution $F$, we have different optimal stopping problems.
-If nothing is known, we have essentially the secretary problem [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989].
+If nothing is known, we have essentially the secretary problem [21, 29].
 [^5] The classical secretary problem considers random order without prior.
-Nonetheless, correaProphetInequalitiesIID2019 showed that the optimal competitive ratio is still $\frac{1}{\mathrm{e}}$ for an unknown prior.
+Nonetheless, Correa et al.
+[14] showed that the optimal competitive ratio is still $\nicefrac{{1}}{{\mathrm{e}}}$ for an unknown prior.
 With complete information of $F$, we have the i.i.d.
-prophet inequality [correaUnknownIIDProphets2021].
+prophet inequality [13].
 This paper considers a new model between the two extremes, which we dub optimal stopping with a predicted prior.
@@
 Hence, upon receiving each value $x_{i}$, the algorithm decides whether to accept based on the number of values $n$, the predicted prior $\tilde{F}$, and the realizations of the values observed thus far including $x_{i}$.
-Following the literature of algorithms with predictions [mitzenmacher2022algorithms], we consider the consistency and robustness of stopping algorithms.
+Following the literature of algorithms with predictions [46], we define the consistency and robustness of stopping algorithms.
 Definition 1 (Consistency).
@@
 2.3 Simplifying Assumptions for Algorithm Analyses
-For our algorithmic results, we adopt two standard assumptions from the literature that simplify the analysis and are without loss of generality.
-Continuous-time Random-order Model.
-Besides the above classical model, which we will refer to as the discrete-time model, we may artificially introduce a random order of arrivals by letting each value $x_{i}$ with $i\in[n]$ arrive in a continuous time horizon $[0,1]$, drawing its arrival time $t_{i}$ independently and uniformly from $[0,1]$.
-Since the values are drawn i.i.d.
-from $F$, the problem remains the same.
-An algorithm for the discrete-time model also works in the continuous-time model by ignoring the arrival time.
-The converse is also true.
-We can simulate the continuous-time model by first drawing $n$ arrival times independently and uniformly from $[0,1]$, and then mapping these arrival times in ascending order to the values in their arrival order.
-Full-support Continuous Priors.
-For the analysis, we further assume that both $F$ and $\tilde{F}$ are continuous with full support on $0,\infty)$.
-This assumption is without loss of generality, e.g., by independently perturbing each value with non-negative noise drawn from an exponential distribution with parameter $\varepsilon\to 0$; see correaProphetSecretaryBlind2021 for a similar treatment.
-For any full-support continuous distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
+For our algorithmic results, we adopt three standard assumptions from the literature that simplify the analysis and are without loss of generality.
+Continuous-time Model.
+Besides the above classical model, which we will refer to as the discrete-time model, we also consider an equivalent continuous-time model.
+In this variant, each value $x_{i}$ is further assigned an arrival time $t_{i}$ drawn independently and uniformly from $[0,1]$, at which time the algorithm observes the value.
+The induced order of arrivals is a uniformly random permutation of the $n$ values, so the discrete-time and continuous-time formulations are equivalent.
+Random Tie-breaking.
+We assume the predicted prior $\tilde{F}$ is continuous. This assumption is without loss of generality for our algorithm, since we can apply the random tie-breaking rule by Correa et al. [17].
+Full-support Continuous Prior.
+For the analysis, we further assume that the prior $F$ is continuous with full support on $0,\infty)$. This assumption is without loss of generality. To see this, we construct a perturbed distribution
+[…]
+Then $F_{\varepsilon}$ is continuous with full support.
+Moreover, with probability at most $1-(1-\varepsilon)^{n}\leq n\varepsilon$, the instance observed by the algorithm will be changed.
+Thus, for any finite $n\in\mathbb{N}$, taking $n\varepsilon\to 0$, the performance is preserved.
+For any distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
 $F^{-1}(q)\coloneqq\inf\big\{x\geq 0:F(x)\geq q\big\},\quad\forall q\in[0,1].$
@@
 Definition 4.
-A value $x$ arriving at time $t$ is best-so-far if $x$ is greater than or equal to the prefix maximum for time $t$.
+A value $x$ arriving at time $t$ is called a record if $x$ is greater than or equal to the prefix maximum for time $t$.
 We consider the following family of algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$.
@@
 Bi-criteria Algorithm $\mathcal{A}(\tilde{F},\theta)$ For each value $x_{i}$ arriving at time $t_{i}$ when no value has been accepted, accept $x_{i}$ if: 1.
-Value $x_{i}$ is best-so-far; and 2.
+Value $x_{i}$ is a record; and 2.
 The cdf value of $x_{i}$ w.r.t.
@@
 The second prior-dependent criterion represents how the algorithm uses the predicted prior to guide its decisions.
-Specifically, if $\theta(t)=1$, the algorithm accepts values with probability 0 at time $t$, regardless of how large the values are.
-If $\theta(t)=0$, the algorithm accepts any value that is best-so-far at time $t$, regardless of the value’s relative standing w.r.t.
+Specifically, if $\theta(t)=1$, the algorithm does not accept any values at time $t$, regardless of how large the values are.
+If $\theta(t)=0$, the algorithm accepts any record at time $t$, regardless of the value’s relative standing w.r.t.
 the predicted prior.
@@
 Proposition 4.
-The algorithm $\mathcal{A}$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
+The algorithm $\mathcal{A}(\tilde{F},\theta)$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
 Proof.
 If $\tilde{F}(y)\leq\theta(s)$, then all values $x$ arriving before time $s$ are rejected because the cdf value is below the threshold, as $\tilde{F}(x)\leq\tilde{F}(y)\leq\theta(s)\leq\theta(t)$ for any $t\leq s$.
-Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be best-so-far.
-If $\tilde{F}(y)>\theta(s)$, then $y$ is best-so-far and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
+Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be a record.
+If $\tilde{F}(y)>\theta(s)$, then $y$ is a record and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
 Corollary 5.
@@
 This family encompasses many classical algorithms as special cases. We present three examples.
-Example 1 (Dynkin’s algorithm [dynkinOptimumChoiceInstant1963], continuous-time variant).
-Dynkin’s algorithm accepts the first best-so-far value after rejecting the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values.
-The continuous-time variant (e.g., feldmanImprovedCompetitiveRatios2011) accepts the first best-so-far value after time $\frac{1}{\mathrm{e}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\frac{1}{\mathrm{e}}]$, and $\theta(t)=0$ for $t\in(\frac{1}{\mathrm{e}},1]$.
-Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-competitive for both MaxProb and MaxExp, without any information about the prior [dynkinOptimumChoiceInstant1963, feldmanImprovedCompetitiveRatios2011].
-Hence, it is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
-Example 2 (Gilbert-Mosteller algorithm [gilbertRecognizingMaximumSequence], continuous-time variant).
-Gilbert and Mosteller calculated the optimal thresholds for MaxProb through backward induction with a correct prior, i.e., $\tilde{F}=F$.
-Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to samuels1982: [^5] The competitive ratio presented in samuels1982 is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
+Example 1 (Dynkin’s algorithm [21], continuous-time variant).
+Dynkin’s algorithm accepts the first record after rejecting the first $\lfloor\nicefrac{{n}}{{\mathrm{e}}}\rfloor$ values.
+The continuous-time variant (e.g., Feldman et al.
+[27]) accepts the first record after time $\nicefrac{{1}}{{\mathrm{e}}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\nicefrac{{1}}{{\mathrm{e}}}]$, and $\theta(t)=0$ for $t\in(\nicefrac{{1}}{{\mathrm{e}}},1]$.
+Dynkin’s algorithm is $\nicefrac{{1}}{{\mathrm{e}}}$-competitive for both MaxProb and MaxExp, without using any information about the prior [21, 27].
+Hence, it is $\nicefrac{{1}}{{\mathrm{e}}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust.
+Example 2 (Gilbert-Mosteller algorithm [33], continuous-time variant).
+With a correct prior $\tilde{F}=F$, the optimal algorithm for the MaxProb objective takes the bi-criteria form according to backward induction.
+In this setting, Gilbert and Mosteller calculated the optimal thresholds in the discrete-time model.
+Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to Samuels [52]: [^6] The competitive ratio presented in Samuels [52] is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
 $\int_{0}^{1}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
 where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
-The continuous-time variant under our framework corresponds to the threshold function given by the following equation:
-$\int_{t}^{1}\frac{(1-s+s\theta(t))^{n-1}-\theta(t)^{n-1}}{1-s}\,\mathrm{d}s=\theta(t)^{n-1}~.$
-It achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm (as a special case of Theorem 2).
+Using a similar backward induction argument, we compute in Lemma 8 the optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ under our bi-criteria framework for any $n\in\mathbb{N}$.
+For each $t$, $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution $q$ of:
+$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\mathrm{d}s=q^{n-1}~.$
+This construction achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm, and it is exactly the form used in the proof of Theorem 2.
 However, the algorithm is merely 0-robust.
@@
 the predicted prior.
-Example 3 (Single-threshold algorithm [ehsaniProphetSecretaryCombinatorial2018], i.i.d. case).
+Example 3 (Single-threshold algorithm [22], i.i.d. case).
 Let $\theta(t)=1-\tfrac{1}{n}$ for $t\in[0,1]$.
-[^6] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\frac{1}{\mathrm{e}}$, which corresponds to $F(x_{i})\approx 1-\frac{1}{n}$ for each $i$ in the i.i.d.
+[^7] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\nicefrac{{1}}{{\mathrm{e}}}$, which corresponds to $F(x_{i})\approx 1-\nicefrac{{1}}{{n}}$ for each $i$ in the i.i.d.
 case.
-The single-threshold algorithm achieves a competitive ratio of $1-\frac{1}{\mathrm{e}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp.
-Hence, it is $(1-\frac{1}{\mathrm{e}})$-consistent for MaxExp.
+The single-threshold algorithm achieves a competitive ratio of $1-\nicefrac{{1}}{{\mathrm{e}}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp.
+Hence, it is $(1-\nicefrac{{1}}{{\mathrm{e}}})$-consistent for MaxExp.
 However, it suffers from the same counterexample and is only 0-robust like Gilbert-Mosteller algorithm.
@@
 Then, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta(t)$ can be any non-increasing function for $t$ between $\lambda_{1}$ and $\lambda_{2}$, while having value $1$ for $0\leq t<\lambda_{1}$ and value 0 for $\lambda_{2}<t\leq 1$.
-We will further show in Section 5 that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$.
-They correspond to a variant of Dynkin’s algorithm that starts to accept best-so-far values at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\frac{1}{\mathrm{e}}$.
-The competitive ratio $\lambda\ln\lambda$ of this algorithm (e.g., feldmanImprovedCompetitiveRatios2011) motivates the following definition of robust threshold functions.
+In Section 5, we will further show that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$.
+They correspond to a variant of Dynkin’s algorithm that starts to accept records at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\nicefrac{{1}}{{\mathrm{e}}}$, whose competitive ratio is $\lambda\ln\lambda$ for both MaxProb and MaxExp (e.g., by Feldman, Naor, and Schwartz [27]).
+This yields the following robustness guarantee, whose proof is deferred to Section 5.
+Lemma 7.
+For any predicted prior $\tilde{F}$, any threshold function $\theta$, denote $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is
+$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\text{-robust}$
+for both MaxProb and MaxExp.
+Motivated by Lemma 7, we now define the robust counterpart of any threshold function.
 Definition 5.
-Consider any feasible robustness parameter $0\leq\beta\leq\frac{1}{\mathrm{e}}$. Let $0\leq\lambda_{1}\leq\frac{1}{\mathrm{e}}\leq\lambda_{2}\leq 1$ be the solutions to:
+For any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, for any threshold function $\theta$, define its corresponding robust threshold function as:
+$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t\leq\lambda_{2}~;\\ 0&\lambda_{2}<t\leq 1~.\end{cases}$
+The robustness guarantee in Lemma 7 depends only on the pair $(\lambda_{1},\lambda_{2})$, which allows us to parameterize robustness directly.
+Consider any feasible robustness parameter $\beta\in[0,\nicefrac{{1}}{{e}}]$.
+Let $0\leq\lambda_{1}\leq\nicefrac{{1}}{{\mathrm{e}}}\leq\lambda_{2}\leq 1$ be the solutions to:
 $\lambda\ln\lambda~=~\beta~.$
+Lemma 7 above immediately implies that any algorithm with the form $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is $\beta$-robust for both MaxProb and MaxExp.
 Readers familiar with the Lambert $\mathcal{W}$ function may note that:
 $\lambda_{1}\coloneqq\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}~,\quad\qquad\lambda_{2}\coloneqq\mathrm{e}^{\mathcal{W}_{0}(-\beta)}~.$ (2)
-For any threshold function $\theta$, define a corresponding robust threshold function as:
-$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t<\lambda_{2}~;\\ 0&\lambda_{2}\leq t\leq 1~.\end{cases}$
-For example, Figure 2 shows the robust counterpart of the threshold function from Gilbert-Mosteller algorithm.
-Figure 2: Illustration of the robust threshold function corresponding to Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{0}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.538$.
+As an illustration, Figure 2 shows the robust counterpart of the threshold function from the Gilbert-Mosteller algorithm with robustness $\beta=\nicefrac{{1}}{{3}}$.
+Figure 2: Illustration of the robust threshold function corresponding to the Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathcal{W}_{-1}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathcal{W}_{0}(-\nicefrac{{1}}{{3}})}\approx 0.538$.
+3.3 Optimal Robust Threshold Functions
+By Lemma 7, once $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$ is fixed, every robust threshold of the form $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is guaranteed to be $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp.
+Therefore, under this certified robustness lower bound, we can optimize consistency by freely choosing the threshold values on $(\lambda_{1},\lambda_{2})$.
+Our main result in this subsection is that, for any $n$, there exist objective-specific optimal thresholds $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ (MaxExp) and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ (MaxProb), whose robust counterparts $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ maximize consistency for every $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$.
+Lemma 8 (Uniqueness and simultaneous optimality of threshold functions).
+For any $n$, any prior distribution $F$ and any objective, either MaxExp or MaxProb, there exists a unique non-increasing threshold function, denoted by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ for MaxExp and by $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ for MaxProb, such that, for any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, the corresponding bi-criteria algorithm induced by the robust threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxExp) or $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxProb) achieves the maximum consistency among all algorithms $\mathcal{A}(F,\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$, where $\theta$ ranges over all threshold functions.
+Here, the function $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is precisely the threshold function used in Example 2.
+We now sketch the intuition behind why the bi-criteria algorithm admits simultaneously optimal threshold functions.
+The proof of Lemma 8 relies on the backward-induction equations and on the existence, uniqueness, and monotonicity of the corresponding optimal thresholds for MaxExp and MaxProb; these technical ingredients are developed in Appendix A.
+Proof Sketch.
+For a fixed objective and prior $F$, consider the dynamic programming formulation of the bi-criteria algorithm with a (not yet specified) threshold function $\theta$.
+At each time $t$, the algorithm compares the immediate payoff from accepting the current value $x$ with the continuation payoff from rejecting and following the same threshold rule in the future.
+The optimal threshold at time $t$ is characterized by an indifference condition between these two options.
+More precisely, suppose a value $x$ arrives at time $t$ and is a record. If $x$ is right at the threshold, i.e., $F(x)=\theta(t)$, then the algorithm rejects $x$, and we have the following.
+  • Prior-free criterion. All future values smaller than $x$ will be rejected, because they cannot be a record.
+  • Prior-dependent criterion. Since $\theta$ is non-increasing in $t$ and $\theta(t)=F(x)$, the first future value larger than $x$ will both exceed the threshold and be a new record.
+In sum, the algorithm will accept the first future value that exceeds $x$.
+The continuation payoff is therefore the expected value (for MaxExp) or success probability (for MaxProb) associated with that future acceptance, which depends only on $t$, $x$, and the prior $F$, and does not depend on the robustness parameters $(\lambda_{1},\lambda_{2})$.
+In Appendix A, we derive that for each objective and each time $0\leq t\leq 1$, the indifference (Bellman) equation that characterizes the optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$ for MaxExp and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ for MaxProb.
+There we show that each equation has a unique solution for every $t$, and that the resulting threshold function is non-increasing in $t$.
+The detailed arguments appear in Lemmas 21 and 24 in the appendix.
+Since the continuation payoffs are independent of $(\lambda_{1},\lambda_{2})$ when the predicted prior is correct, the same objective-specific threshold solves the indifference equation of time $t$ for every choice of $\lambda_{1}\leq t\leq\lambda_{2}$, and for any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, the robust threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxExp) or $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxProb) maximizes consistency among all $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ in the bi-criteria framework.
+This concludes our proof. ∎
+For the MaxProb objective, setting the robustness requirement to $\beta=0$ makes the predicted-prior problem coincide with the classical i.i.d.
+prophet inequality with the MaxProb objective.
+By Lemma 8, the threshold $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is the unique consistency-maximizing threshold for this objective within our continuous-time framework.
+As promised in Example 2, we show that it achieves the optimal $0.580$ competitive ratio in Section 7, where we prove Theorem 2.
+In contrast, for MaxExp, our optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ is not directly comparable to those of optimal i.i.d.
+prophet inequality algorithms, which rely solely on a threshold-based acceptance criterion without any record-based condition.
+It remains open whether one can attain the $0.745$ worst-case consistency guarantee within this structural class, or whether any such record-based threshold policy must fall strictly below this benchmark.
 3.4 Brief Discussion on Algorithms Outside the Framework
-In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is best-so-far.
+In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is a record.
 With a correct prior, the optimal policy by backward induction takes this form, where the threshold is the expected accepted value if the algorithm rejects the current value.
-For example, see the work by krengel1977semiamarts on general prophet inequality and the more recent research [correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021, harbNewProphetInequalities2025] for i.i.d.
+For example, see the work by Krengel and Sucheston [42] on general prophet inequality and the more recent research [16, 44, 37] for i.i.d.
 prophet inequality.
-These algorithms are outside our framework, where the taken value must be best-so-far, in addition to surpassing the time-dependent threshold.
-We stress that the best-so-far requirement is critical for robustness.
-The mentioned algorithms that have no best-so-far requirements are only 0-robust.
+These algorithms are outside our framework, where the taken value must be a record, in addition to surpassing the time-dependent threshold.
+We stress that record-based requirement is critical for robustness.
+The mentioned algorithms without record-based requirements are only 0-robust.
 To see this, consider a predicted prior supported on values much smaller than those in the real prior.
-Then, all values from the real prior will have cdf value essentially $1$ w.r.t.
+Then, all values from the real prior will have a cdf value essentially $1$ w.r.t.
 the predicted prior.
-As a result, the algorithm accepts the first value arrived when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
+As a result, the algorithm accepts the first value that arrives when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
 Meanwhile, our framework is limited in that it does not adjust thresholds adaptively based on the observed values.
-Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [dynkinOptimumChoiceInstant1963] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.5801$-competitive algorithm [gilbertRecognizingMaximumSequence] in the MaxProb setting or the $0.745$-competitive algorithm [correaProphetInequalitiesIID2019] in the MaxExp setting if the observed values are consistent with the predicted prior.
+Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [21] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.580$-competitive algorithm [33] in the MaxProb setting or the $0.745$-competitive algorithm [14] in the MaxExp setting if the observed values are consistent with the predicted prior.
 We consider a static threshold function to make the analyses tractable.
@@
 Theorem 9.
-For any $n,k\in\mathbb{N}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
-By our convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$.
+For any $n,k\in\mathbb{N}$, for any predicted prior $\tilde{F}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
+By convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$.
 Importantly, the maximum of $k$ values drawn i.i.d.
 from $\tilde{F}^{\nicefrac{{1}}{{k}}}$ follows distribution $\tilde{F}$.
-We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values drawn i.i.d.
+We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values $x_{1},x_{2},\ldots,x_{n}$ to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values $x^{\prime}_{11},\ldots,x^{\prime}_{1k},\ldots,x^{\prime}_{n1},\ldots,x^{\prime}_{nk}$ drawn i.i.d.
 from $\tilde{F}^{\nicefrac{{1}}{{k}}}$.
-Let $x_{1},\dots,x_{n}$ be the $n$ values by their arrival order in the original instance.
-Further, consider an instance of $nk$ values $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ drawn i.i.d.
-from distribution $F^{\nicefrac{{1}}{{k}}}$ with sorted arrival times $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ drawn i.i.d.
-and uniformly from $[0,1]$.
-Definition 6 (Coupling for Sharding).
-Let $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$, and let $s_{i}$ be the arrival time of $\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}$.
-We use $s_{i}$ instead of $t_{i}$ to avoid confusion, as $s_{i}$ is not uniform over $[0,1]$.
 Naïvely, we would like to simulate algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ by sampling $x^{\prime}_{ij}$, $j\in[k]$, from $F^{\nicefrac{{1}}{{k}}}$ conditioned on the realization of $x_{i}$, and accept value $x_{i}$ if and only if one of the $x^{\prime}_{ij}$’s is accepted by $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
-This is the sharding technique by harbNewProphetInequalities2025 for the prophet inequality setting.
+This is the sharding technique by Harb [37] for the prophet inequality setting.
 Unfortunately, the algorithm does not know $F$ in the problem of optimal stopping with a predicted prior, and thus cannot sample $x^{\prime}_{ij}$ from $F^{\nicefrac{{1}}{{k}}}$.
@@
 We will show that this partial simulation is at least as good as the full simulation in the original sharding method.
-Further, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis.
+In contrast to the sharding technique, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis.
+Our algorithm is defined as follows.
 Bi-criteria Algorithm with Implicit Sharding $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ 1. Sample nknk arrival times i.i.d. from $\mathrm{Unif}[0,1]$ and sort them as:
@@
 3.
-For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if: (a) Value $x_{i}$ is best-so-far; and (b) The cdf value of $x_{i}$ w.r.t.
-$\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if:
+For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if (a) Value $x_{i}$ is a record; and (b) The cdf value of $x_{i}$ w.r.t.
+$\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if
               $\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})~.$
+Next, we design a coupling that gives a joint distribution preserving the marginal distributions of the random variables used in our algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ and in the simulated algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, respectively.
+Definition 6 (Coupling).
+Sample $nk$ i.i.d.
+values from $F^{\nicefrac{{1}}{{k}}}$.
+For each value, sample an arrival time i.i.d.
+from $\mathrm{Unif}[0,1]$.
+Then, we index the values by the ascending order of their arrival times.
+Denote the values by $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ and the sorted times by
+$t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}.$
+For each $i$, define
+$x_{i}\coloneqq\max_{j\in[k]}x^{\prime}_{ij},\qquad s_{i}\coloneqq t^{\prime}_{ij_{i}},\quad\text{where }j_{i}\in\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}.$
+Under this coupling, each $x_{i}$ has distribution $F$, and $\max_{i\in[n]}x_{i}=\max_{i\in[n],j\in[k]}x^{\prime}_{ij}$ pointwise.
+Also, because the shards in each block are i.i.d., the argmax index $j_{i}$ is uniform over $[k]$, so $s_{i}$ has the same distribution as Step 2 in $\mathcal{A}_{\mathrm{sharding}}$.
+We stress that the algorithm does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s. They are only used in the coupling for analysis.
 Proof of Theorem 9.
-It suffices to show that for any real prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
+We establish the pointwise value dominance property: for any prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
 Suppose algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ accepts value $x_{i}$, and $x^{\prime}_{ij}=x_{i}$ is the corresponding value in the coupled instance for algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
-Since $x_{i}$ is best-so-far, $x^{\prime}_{ij}$ is also best-so-far because
-$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}\leq i,\;j^{\prime}\in[k].$
+Since $x_{i}$ is a record, $x^{\prime}_{ij}$ is also a record because
+$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}},t^{\prime}_{ij}=s_{i}\geq t^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}<i,\;j^{\prime}\in[k]\text{ or }i^{\prime}=i,\;j^{\prime}\leq j.$
 Further, the cdf value of $x^{\prime}_{ij}$ is greater than the threshold at time $t^{\prime}_{ij}$, as
 $\tilde{F}^{\nicefrac{{1}}{{k}}}(x^{\prime}_{ij})=\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})=\theta(t^{\prime}_{ij})~.$
-Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is best-so-far.
-Taking expectations over the coupling completes the proof.
-∎
+Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is a record.
+By this pointwise value dominance, we obtain the following conclusion for each objective:
+  1. 1.
+    For MaxExp: The conclusion follows directly by taking expectations over the coupling.
+  2. 2.
+    For MaxProb: The conclusion follows directly by taking expectations over all points in the coupling where $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts the maximum value.
+This completes the proof. ∎
 5 Proof of Robustness
-We analyze the robustness of our algorithm for MaxProb (which then implies robustness for MaxExp) in a stronger setting that fixes the realized values and randomizes only the arrivals.
-This problem is also known as the game of Googol [fergusonWhoSolvedSecretary1989, gnedin1994solution].
+In Section 3 we stated Lemma 7, which bounds the robustness of $\mathcal{A}(\tilde{F},\theta)$ in our stochastic model.
+We prove it in this section.
+The argument proceeds via a stronger “fixed values, random arrivals” model, also known as the game of Googol [29, 34], in which the values $x_{1},x_{2},\ldots,x_{n}$ are treated as deterministic and only the arrival order is random.
+The key step is the following bound in the game-of-Googol setting.
 Proposition 10.
-For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value with probability at least
+For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ accepts the maximum value with probability at least
 $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}~,$
 over the random arrivals of the values.
-This implies the robustness guarantees for both MaxProb and MaxExp as corollaries through a standard argument.
-Corollary 9.
-For any predicted prior $\tilde{F}$, any threshold function $\theta$, and any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp.
-Proof.
+We now derive Lemma 7 from Proposition 10.
+Proof of Lemma 7.
+Fix a predicted prior $\tilde{F}$ and a threshold function $\theta$.
+By definition, for $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$, we have $\theta=\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$.
+Thus, it suffices to show that $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust.
 By the law of total probability, we have
-$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},\dots,x_{n}}\Big[\,\operatorname{\mathbf{Pr}}_{t_{1},\dots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},\dots,x_{n}\,\big]\,\Big]~.$
-By Proposition 8, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity.
+$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},x_{2},\ldots,x_{n}}\big[\operatorname{\mathbf{Pr}}_{t_{1},t_{2},\ldots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},x_{2},\ldots,x_{n}\,\big]\big]~.$
+Applying Proposition 10, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity.
 This proves robustness for MaxProb.
-Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, i.e., the same robustness for MaxExp.
+Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, since for any realization of the values, the accepted value equals the maximum value with the same probability as in the MaxProb case.
+This proves robustness for MaxExp as well.
 ∎
@@
 Lemma 11.
-For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
+For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
 $\int_{0}^{1}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{n})>\theta(t)\}}}\bigg((1-t)^{n-1}+\sum_{i=1}^{n-1}(1-t)^{n-1-i}\int_{0}^{t}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(s)\}}}\,\mathrm{d}s\bigg)\,\mathrm{d}t~.$ (3)
 Proof.
-Since $x_{n}$ is always best-so-far, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$.
+Since $x_{n}$ is always a record, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$.
 By Proposition 4, the latter is equivalent to the event that the prefix maximum for time $t_{n}$ has cdf under $\tilde{F}$ at most the threshold at its arrival time.
@@
 First, the prefix maximum is 0 if no value arrives before $t_{n}$, which occurs with probability $(1-t_{n})^{n-1}$.
-Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$.
+Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},x_{i+2},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$.
 In this case, we also require $\tilde{F}(x_{i})\leq\theta(t_{i})$ for the arrival time $t_{i}\in[0,t_{n})$, contributing
 $(1-t_{n})^{n-1-i}\int_{0}^{t_{n}}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(t_{i})\}}}\,\mathrm{d}t_{i}~.$
-Putting together and changing variables $t_{n}$ with $t$ and $t_{i}$ with $s$, the lemma follows. ∎
+Putting everything together and changing variables $t_{n}$ to $t$ and $t_{i}$ to $s$, the lemma follows. ∎
 We are now ready to prove Proposition 10.
@@
 First, we fix $\theta$ and characterize the worst-case predicted prior $\tilde{F}$.
-Equation 2 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\dots\leq\tilde{F}(x_{n})\leq 1$.
-Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n-1})$.
+Equation 3 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\tilde{F}(x_{2})\leq\cdots\leq\tilde{F}(x_{n})\leq 1$.
+Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\tilde{F}(x_{2}),\ldots,\tilde{F}(x_{n-1})$.
 Therefore, given any fixed threshold function $\theta$, Equation 3 is minimized when all $\tilde{F}(x_{i})$ are equal to some $q\in[0,1]$.
 In this case, the threshold function is effectively binary.
-A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n})$ all surpass the threshold.
+A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\tilde{F}(x_{2}),\ldots,\tilde{F}(x_{n})$ all surpass the threshold.
 Similarly, a threshold $\theta(t)\geq q$ is equivalent to $\theta(t)=1$, as none surpasses it.
@@
 $\theta(t)=\begin{cases}1,&\text{if }t\leq\lambda~;\\ 0,&\text{otherwise~.}\end{cases}$
-This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\frac{1}{\mathrm{e}}$.
-This is $(-\lambda\ln\lambda)$-competitive (e.g., feldmanImprovedCompetitiveRatios2011).
+This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\nicefrac{{1}}{{\mathrm{e}}}$.
+This algorithm is $(-\lambda\ln\lambda)$-competitive (e.g., Feldman et al.
+[27]).
 Since $\lambda\ln\lambda$ is concave in $\lambda$ on $[0,1]$, its minimum over $\lambda_{1},\lambda_{2}]$ is attained at an endpoint, yielding the bound $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$.
@@
 6 Algorithm for MaxExp: Proof of Theorem 1
-By the implicit sharding method (Theorem 7), it suffices to design an algorithm $\mathcal{A}(\tilde{F},\theta)$ under the bi-criteria framework, and prove the desired consistency and robustness guarantees as the number of values $n$ tends to infinity.
-Further, by Corollary 9, the robustness guarantee holds for any robust threshold $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$.
-The rest of the section focuses on the consistency guarantee, for any fixed robustness parameter $\beta\in[0,\frac{1}{e}]$ and the corresponding $\lambda_{1},\lambda_{2}$.
-Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will only use $F$ throughout the section.
-To formally study the limit as $n\to\infty$, we will design the threshold function for the extremal distribution $F^{n}$.
-More precisely, we will design a threshold function $\theta$ and analyze the consistency of algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ for $n$ values, and in particular, the limit when $n$ tends to infinity.
-As a standard approach, we design the algorithm so that for any $\ell\geq 0$, it holds
-$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~.$ (3)
-This implies the consistency guarantee as
+In this section we prove Theorem 1.
+We fix a $\beta\in[0,\nicefrac{{1}}{{e}}]$, and fix $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$ as the two roots of $\lambda\ln\lambda=\beta$.
+Once there exists a non-increasing function $\bar{\theta}\colon[\lambda_{1},\lambda_{2}]\to[0,1]$ that satisfies Equation (1), with a little abuse of notation, we extend the definition of $\bar{\theta}$ to the entire interval $[0,1]$ by setting $\bar{\theta}(t)=1$ for $t<\lambda_{1}$ and $\bar{\theta}(t)=0$ for $t>\lambda_{2}$.
+By Theorem 9, it suffices to construct an algorithm that is both $\alpha$-consistent and $\beta$-robust as the number of values $n$ tends to infinity.
+We show that $\mathcal{A}(\tilde{F},\bar{\theta}^{\nicefrac{{1}}{{n}}})$ satisfies these guarantees.
+The $\beta$-robustness directly follows from Lemma 7, since $\bar{\theta}^{\nicefrac{{1}}{{n}}}=\bar{\theta}^{\nicefrac{{1}}{{n}}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ by the extension above.
+The rest of the section focuses on the consistency guarantee.
+Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will write it as $F$ throughout the section.
+The remaining task is to show that $\mathcal{A}(F,\bar{\theta}^{\nicefrac{{1}}{{n}}})$ is $\alpha$-competitive as $n\to\infty$.
+As a standard approach in competitive analysis, we show that for any $\ell\geq 0$, it holds
+$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~,$ (4)
+i.e., $x_{\textsc{Alg}}$ stochastically dominates $\alpha\cdot x_{\textsc{Opt}}$. This implies the desired consistency
 $\operatorname{\mathbf{E}}\big[\,x_{\textsc{Alg}}\,\big]~=~\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]\,\mathrm{d}\ell~\geq~\alpha\cdot\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]\,\mathrm{d}\ell~=~\alpha\cdot\operatorname{\mathbf{E}}\big[\,x_{\textsc{Opt}}\,\big]~.$
+Remark 1.
+The algorithm designed in this section does not use the theoretically optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ derived in Section 3.3; that threshold maximizes consistency among all algorithms with robust thresholds of the form $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$.
+For the MaxExp objective, $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ is given by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$, where $x^{*}(t)$ is the unique solution of the Bellman equation (see Equation 17), so $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ depends on the prior distribution $F$ through quantities such as the conditional expectation $\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]$ and the quantile $F(x)$.
+As a result, the consistency guarantees $\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}\geq\ell]$ under $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ are expressed in terms of these prior-dependent objects.
+In this section we instead work with a suboptimal but analytically more tractable threshold that is independent of the prior $F$, and we analyze its consistency relative to $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ for each fixed prior.
+The proof consists of two parts. First, we simplify Equation 4 to give a sufficient condition that $\bar{\theta}$ needs to satisfy. Then, we verify that $\bar{\theta}$ satisfies these conditions.
 6.1 Sufficient Conditions for Infinitely Many Values
-We consider a general threshold function $\theta$ and derive an expression for the left-hand side of Equation 3, i.e., the probability that the algorithm accepts a value that is at least $\ell$.
+First, to simplify the left-hand side of Equation 4, we use the following lemma to calculate the probability a bi-criteria algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}})$ accepts a value that is at least $\ell$.
 Lemma 12.
@@
 Multiplying by $n$ by symmetry across all $x_{i}$’s proves the lemma. ∎
-For any $y\in[0,1]$, if we let $n$ tend to infinity while keeping the same $F(\ell)^{n}=y$, the desired Inequality (3) becomes
-$\int_{y}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
-where we use the fact that $\big(1-t+tz^{\nicefrac{{1}}{{n}}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity.
-[^7] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n$ tends to infinity and consider any $\varepsilon>0$.
+Now, focus on algorithm $\mathcal{A}(F,\bar{\theta}^{\nicefrac{{1}}{{n}}})$. For any $y\in[0,1]$, if we let $n\to\infty$ while keeping the same $F(\ell)^{n}=y$, the desired Inequality (4) becomes
+$\int_{y}^{1}\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
+where we use the fact that $\big(1-t+tz^{\frac{1}{n}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity.
+[^8] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n\to\infty$ and consider any $\varepsilon>0$.
 By the uniform convergence of the stated limit for $q\in[\varepsilon,1]$, we can apply implicit sharding to ensure Equation 4 for $\alpha=\alpha^{*}(1-\varepsilon)$ and any $q=F(\ell)^{n}\in[\varepsilon,1]$.
-Finally, Equation 3 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most a $1-\varepsilon$ factor.
+Finally, Equation 4 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most the $1-\varepsilon$ factor.
 Define the function with support set $[0,1]$ as
-$g(q)~\coloneqq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
-Then, it suffices to design a threshold function $\theta$ such that for any $q\in[0,1]$, it holds
+$g(q)~\coloneqq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\bar{\theta}(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
+Then, it suffices to show that for any $q\in[0,1]$, it holds
 $g(q)~\geq~\alpha~.$
-6.2 Proof of Theorem 1
-Next, we focus on a robust threshold function $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, and for ease of notation we continue to write $\theta$ instead of $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, restricting $\theta$ from now on such that
+6.2 Verifying the Sufficient Condition
+Recall the assumed properties of $\bar{\theta}$ by Equation (1):
   1. 1)
-    $\theta(t)=1$ for $t\in[0,\lambda_{1})$; and
+    $\bar{\theta}$ is left-continuous, which implies $\bar{\theta}(\lambda_{1})=1$, and non-increasing; and
   2. 2)
-    $\theta(t)=0$ for $t\in(\lambda_{2},1]$.
-Further, recall the assumed properties in Theorem 1:
-  1. 3)
-    $\theta$ is left-continuous, which implies $\theta(\lambda_{1})=1$, and non-increasing; and
-  2. 4)
     For any $z\in[\lambda_{1},\lambda_{2}]$, we have
-    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\theta(z)~.$
-Consider any $q$ with $\theta(\lambda_{2})\leq q\leq\theta(\lambda_{1})=1$. Let $z=\theta^{-1}(q)$, and then by property 4), we have
-$\alpha$ $~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\theta(\max\{s,\theta^{-1}(q)\})^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
-$~=~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),\theta(\theta^{-1}(q))\big\}^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t.$
-By left-continuity of $\theta$ and the definition of $\theta^{-1}$, we have $\theta(\theta^{-1}(q))\geq q$. Therefore,
-$\alpha~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\theta(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q).$
-Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\theta(\lambda_{2}))$ by reducing it to the first case.
+    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\bar{\theta}(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\bar{\theta}(z)~.$
+Consider any $q$ with $\bar{\theta}(\lambda_{2})\leq q\leq\bar{\theta}(\lambda_{1})=1$. Let $z=\bar{\theta}^{-1}(q)$, and then by property 2), we have
+$\alpha$ $~\leq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\bar{\theta}(\max\{s,\bar{\theta}^{-1}(q)\})^{t}}{\bar{\theta}(\bar{\theta}^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
+$~=~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),\bar{\theta}(\bar{\theta}^{-1}(q))\}^{t}}{\bar{\theta}(\bar{\theta}^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t~.$
+By left-continuity of $\bar{\theta}$ and the definition of $\bar{\theta}^{-1}$, we have $\bar{\theta}(\bar{\theta}^{-1}(q))\geq q$. Therefore,
+$\alpha~\leq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q)~.$
+Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\bar{\theta}(\lambda_{2}))$ by reducing it to the first case.
 Lemma 13.
-For any $q\in[0,\theta(\lambda_{2})]$, we have
-$g(q)\geq g\big(\theta(\lambda_{2})\big).$
+For any $q\in[0,\bar{\theta}(\lambda_{2})]$, we have
+$g(q)\geq g\big(\bar{\theta}(\lambda_{2})\big)~.$
 Proof.
-For any $q$ in the range $0,\theta(\lambda_{2})]$, we have $\theta^{-1}(q)=\lambda_{2}$.
-Also, observe that $\theta(s)\geq\theta(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\theta(s)=0$ for $s\in[\lambda_{2},1]$.
+For any $q$ in the range $0,\bar{\theta}(\lambda_{2})]$, we have $\bar{\theta}^{-1}(q)=\lambda_{2}$.
+Also, observe that $\bar{\theta}(s)\geq\bar{\theta}(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\bar{\theta}(s)=0$ for $s\in[\lambda_{2},1]$.
 Hence,
 $g(q)=\int_{\lambda_{2}}^{1}\int_{0}^{\lambda_{2}}\frac{1}{t}q^{-(1-t)}\,\mathrm{d}s\,\mathrm{d}t=\int_{\lambda_{2}}^{1}\frac{\lambda_{2}}{t}q^{-(1-t)}\,\mathrm{d}t~,$
-which is decreasing in $q$. ∎
+which is decreasing in $q$, implying the desired inequality. ∎
 7 Algorithm for MaxProb: Proof of Theorem 2
-We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value when the predicted prior is correct.
+This section proves Theorem 2 by establishing the optimal consistency-robustness trade-off for the MaxProb objective.
+Recall that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is the threshold function used by the continuous-time Gilbert-Mosteller algorithm (Example 2), where for each $t\in[0,1]$, the value $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution $q$ to
+$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\mathrm{d}s=q^{n-1}~.$ (5)
+By the Implicit Sharding technique (Theorem 9), for any $\beta\in[0,\nicefrac{{1}}{{e}}]$, we can focus on the trade-off of the bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ in the limiting case when $n\to\infty$, where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
+Then, we show that this trade-off is tight within the bi-criteria framework by showing that no algorithm in the form of $\mathcal{A}(\tilde{F},\theta)$ can achieve a better consistency-robustness trade-off.
+7.1 Asymptotic Performance Analysis
+With Lemma 7, the $\beta$-robustness guarantee of $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ follows directly.
+It remains to show that the algorithm is $\alpha$-consistent as $n\to\infty$, where
+$\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\,\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
+To this end, we first derive an explicit expression for the probability that a bi-criteria algorithm accepts the maximum value under a correct prior for finite $n$.
 Lemma 14.
-For any $n\in\mathbb{N}^{+}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ is
-$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s$
-competitive for MaxProb.
+For any $n\in\mathbb{N}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ accepts the maximum value with probability
+$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s\,.$
 Proof.
@@
 $n\int_{0}^{1}\int_{\theta(t)}^{1}\int_{0}^{t}\frac{1}{t}\,\big(\,t\min\{\theta(s),q\}+(1-t)q\,\big)^{n-1}\,\mathrm{d}s\,\mathrm{d}q\,\mathrm{d}t~.$
-Changing the order of integration and simplifying yields the stated expression for $\Gamma_{n}(\theta)$. ∎
+Changing the order of integration and simplifying yield the expression for $\Gamma_{n}(\theta)$. ∎
+With the probability guarantee for general threshold derived in Lemma 14, we next evaluate the consistency achieved by our choice of thresholds for any $\beta\in[0,\nicefrac{{1}}{{\mathrm{e}}}]$.
 Lemma 15.
-For any $n\in\mathbb{N}^{+}$ and any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, define threshold function $\theta_{n}^{*}$ such that for any $s\in[0,1]$, $\theta_{n}^{*}(s)$ is the unique solution to
-$\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t~=~\theta_{n}^{*}(s)^{n-1}~.$ (4)
-We have
-$\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\geq~-\lambda_{1}\ln\lambda_{1}+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
-where equality holds when $n$ tends to infinity.
-Note that $\theta_{n}^{*}$ is exactly the threshold function used in the continuous-time Gilbert-Mosteller algorithm (Example 2). We defer the proofs of some lemmas to Appendix A.
+For any $\beta\in[0,\nicefrac{{1}}{{e}}]$, suppose $\lambda_{1}$, $\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
+For any $n\in\mathbb{N}$, recall that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution to Equation 5 for any $t\in[0,1]$.
+Then
+$\lim_{n\to\infty}\Gamma_{n}\big(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~=~\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
 Proof.
-By the expression for $\Gamma_{n}(\theta)$ in Lemma 13, we derive the definition of $\theta_{n}^{*}$ from the first-order condition for $\theta(t)$, i.e.,
-[…]
-Further, $\theta_{n}^{*}(s)$ is decreasing in $s$ because Equation 4 can be rewritten as
-$\int_{s}^{1}\frac{\big(\tfrac{1-t}{\theta_{n}^{*}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t~=~1~,$ (5)
-except for the boundary condition $\theta_{n}^{*}(1)=0$.
-In sum, $\theta_{n}^{*}$ is a threshold function satisfying the first-order conditions for $t\in[0,1]$.
-Therefore, $\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is a threshold function satisfying the first-order conditions for $t\in[\lambda_{1},\lambda_{2}]$ on one hand, and the restrictions of being $1$ for $t\in[0,\lambda_{1})$ and 0 for $t\in(\lambda_{2},1]$ on the other hand.
-We next write the competitive ratio $\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ as
-$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{*}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
-We now focus on the limit case when $n$ tends to infinity.
-This is without loss of generality, either by the implicit sharding technique, or by the fact that the above ratio is non-increasing in $n$ (see Lemma 19 in Section A.1).
-In particular, the latter means that the stated consistency and robustness guarantees also hold directly for the bi-criteria algorithm without implicit sharding.
-When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}$ by basic calculus, and the last term tends to 0.
-It remains to show that, as $n\to\infty$, the integrand of the second term converges to
+The probability $\Gamma_{n}\big(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ can be written as
+$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n}-t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
+When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}=\beta$ by basic calculus, and the last term tends to 0.
+It remains to show that, as $n\to\infty$, the inner integrand of the second term converges to
 $\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t~.$
-By the definition of $\theta_{n}^{*}$ (Equation 4), the integrand equals
-$\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{*}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t$
-$\qquad={}\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n-1}}{t}\,\mathrm{d}t~.$
-Finally, we prove in Section A.2 (Lemma 20) that
-$\theta_{n}^{*}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$
-uniformly on compact subsets of $[0,1)$, which yields the above limit as $n\to\infty$.
+By the definition of $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ in Equation 5, the integrand equals
+$\quad\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n}-t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)\big)^{n-1}-\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n-1}}{1-t}\,\mathrm{d}t$
+$=\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n-1}}{t}\,\mathrm{d}t~.$
+Finally, we use the asymptotic expansion of the threshold function proven in Appendix C:
+$\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$
+which holds uniformly on compact subsets of $[0,1)$.
+Substituting this limit into the integral yields the lemma.
 [^9] Rigorously, for any $\varepsilon>0$, uniform convergence over $0,1-\varepsilon]$ implies the convergence of $\int_{s}^{1-\varepsilon}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t$.
 The difference from the ideal term is bounded by $\int_{1-\varepsilon}^{1}\frac{1}{t}\,\mathrm{d}t=\mathcal{O}(\varepsilon)$, which vanishes as $\varepsilon\to 0$.
+7.2 Tightness of the Trade-off
+To complete the proof of Theorem 2, we show that no bi-criteria algorithm can achieve a better consistency-robustness trade-off.
+First, we show that for any bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta)$, the robustness guarantee given by Lemma 7 is tight.
+Lemma 16.
+For any $n\in\mathbb{N}$, any threshold function $\theta$, suppose $\lambda_{1}=\sup\{0\leq t\leq 1\colon\theta(t)=1\}$, $\lambda_{2}=\inf\{0\leq t\leq 1\colon\theta(t)=0\}$.
+Then, for any prior $F$ and any $\varepsilon>0$, there exists a predicted prior $\tilde{F}$ such that, when values are drawn i.i.d.
+from $F$, algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts a maximum value with probability at most
+$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}+\varepsilon~.$
+Proof.
+Fix any prior $F$ and $\varepsilon>0$. We prove two bounds, corresponding to “small” and “large” predicted priors.
+Case 1: Small predicted prior.
+Define $\tilde{F}_{\mathrm{small}}(x)=1$ for all $x\geq 0$ (equivalently, a point mass $1$ at 0).
+By monotonicity of $\theta$, for any $t>\lambda_{1}$ we have $\theta(t)<1$, so the second acceptance criterion $\tilde{F}_{\mathrm{small}}(x_{i})>\theta(t_{i})$ always holds.
+Hence, up to a measure-zero boundary at $t=\lambda_{1}$, the algorithm rejects all values before $\lambda_{1}$ and then accepts the first record, i.e., Dynkin’s continuous-time variant with transition point $\lambda_{1}$.
+By the standard analysis of that variant (e.g., Feldman et al.
+[27]), its MaxProb ratio equals $\lambda_{1}\ln\lambda_{1}$.
+Hence
+$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{small}},\theta)=x_{\textsc{Opt}}\big]\leq-\lambda_{1}\ln\lambda_{1}~.$
+Case 2: Large predicted prior.
+Since $F(\infty)=1$, there exists $m\geq 0$ such that
+$F(m)^{n}\geq 1-\varepsilon~.$
+Define
+$\tilde{F}_{\mathrm{large}}(x)=\begin{cases}0,&x<m+1~;\\ 1,&x\geq m+1~.\end{cases}$
+Let
+$E\coloneqq\{X_{1}\leq m,X_{2}\leq m,\ldots,X_{n}\leq m\}~,$
+so $\operatorname{\mathbf{Pr}}[E]\geq 1-\varepsilon$.
+On event $E$, all realized samples satisfy $\tilde{F}_{\mathrm{large}}(X_{i})=0$.
+Therefore, for any $t<\lambda_{2}$, since $\theta(t)>0$, the second acceptance criterion fails and the algorithm cannot accept before $\lambda_{2}$.
+Consequently, conditioned on $E$, this policy cannot outperform the policy that starts accepting records at time $\lambda_{2}$ (Dynkin variant), so
+$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{large}},\theta)=x_{\textsc{Opt}}\,\big|\,E\big]\leq-\lambda_{2}\ln\lambda_{2}~.$
+Consequently,
+$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{large}},\theta)=x_{\textsc{Opt}}\big]\leq\operatorname{\mathbf{Pr}}[E]\cdot(-\lambda_{2}\ln\lambda_{2})+\operatorname{\mathbf{Pr}}[E^{c}]\leq-\lambda_{2}\ln\lambda_{2}+\varepsilon~.$
+Finally, choose $\tilde{F}_{\mathrm{small}}$ if $\lambda_{1}\ln\lambda_{1}\leq-\lambda_{2}\ln\lambda_{2}$, and choose $\tilde{F}_{\mathrm{large}}$ otherwise. In both cases,
+$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]\leq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}+\varepsilon~.$
+If one insists on full-support continuous predicted priors, the same conclusion follows up to an arbitrarily small additive term by the smoothing argument in Section 2. ∎
+We now complete the proof of Theorem 2. Fix any bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta)$, and define
+$\lambda_{1}^{\prime}=\sup\{0\leq t\leq 1\colon\theta(t)=1\}~,\qquad\lambda_{2}^{\prime}=\inf\{0\leq t\leq 1\colon\theta(t)=0\}~,$
+$\beta\coloneqq\min\{-\lambda_{1}^{\prime}\ln\lambda_{1}^{\prime},-\lambda_{2}^{\prime}\ln\lambda_{2}^{\prime}\}~.$
+By Lemma 16, for every $\varepsilon>0$ there exists a misspecification under which this algorithm succeeds with probability at most $\beta+\varepsilon$, hence its robustness is at most $\beta$.
+Now let $\lambda_{1},\lambda_{2}$ be the two roots of $\lambda\ln\lambda=\beta$.
+By the concavity of $\lambda\ln\lambda$ over $[0,1]$, we have $\lambda_{1}^{\prime}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{2}^{\prime}$.
+This means $\theta=\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$.
+By Lemma 8, for these fixed $(\lambda_{1},\lambda_{2})$ and any $n$, the maximum consistency within the class $\mathcal{A}(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is achieved by $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$.
+Therefore, no $\beta$-robust bi-criteria algorithm can have consistency larger than $\Gamma_{n}(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$.
+Taking $n\to\infty$ and applying Lemma 15, this upper bound converges to
+$\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
+Hence, the trade-off in Theorem 2 is tight within the bi-criteria framework.
 8 Hardness for MaxProb
@@
 To derive the linear constraints in polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$, we consider two representations of online algorithms when the predicted prior is $\tilde{F}$ and the real prior is one of the conditional distributions in the above family $\mathcal{F}$.
-The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., buchbinderSecretaryProblemsLinear2014, correaSampledrivenOptimalStopping2021).
+The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., Buchbinder et al.
+[8], Correa et al.
+[12]).
 Definition 8 (Conditional Acceptance Probabilities).
@@
 $\operatorname{\mathbf{Pr}}\big[\textsc{Alg}=t\mid\bm{x}\big]~=~\textsc{Acc}_{t}(\bm{x}_{1:t})\prod_{i\in[t-1]}\big(1-\textsc{Acc}_{i}(\bm{x}_{1:i})\big)~.$
-Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a best-so-far value $x_{t}$ at time $t$ is independent of smaller values $\bm{x}_{1:t-1}$ in the past.
+Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a record $x_{t}$ at time $t$ are independent of smaller values $\bm{x}_{1:t-1}$ in the past.
 Definition 9 (Minor-oblivious Algorithms).
@@
 We formalize this characterization in the following proposition, which may be of independent interest.
-The proof is deferred to Section C.1.
+The proof is deferred to Section D.1.
 Proposition 19.
@@
 $\textsc{Rej}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
-We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ to simplify some expressions in the rest of the section.
-There is a one-to-one correspondence between the two representations (see Section C.2). We state below the direction to be used in the proof of Theorem 15.
+We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq 1$ to simplify some expressions in the rest of the section.
+There is a one-to-one correspondence between the two representations (see Section D.2). We state below the direction to be used in the proof of Theorem 17.
 Lemma 20.
@@
 8.3 Proof of Theorem 17
-Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not best-so-far, since we consider MaxProb.
+Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not a record, since we consider MaxProb.
 We next verify that $y_{t,\ell}=\textsc{Rej}_{t}(\ell)$ is feasible w.r.t.
@@
 By the definition of the minor-oblivious rejection probabilities, and the artificial definition of $\textsc{Rej}_{0}(\ell)$, we obtain Constraints (6) and (7):
-$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,\qquad\forall t\in[n],\;\forall\ell\in[K]~.$
+$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,~~\forall t\in[n],\;\forall\ell\in[K]~.$
 Bounds for Acceptance Probabilities.
@@
 $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{k}\big[\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
-We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not best-so-far.
+We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not a record.
 By the relation between $\tilde{F}=F_{K}$ and $F_{k}$, we can rewrite it as:
@@
 For any $\lambda\in[0,1]$, we solve a linear program with variables $\{y_{t,\ell}\}_{t\in[n],\,\ell\in[K]}\cup\{\alpha,\beta\}$ to maximize:
-$\lambda\cdot\alpha+(1-\lambda)\cdot\beta$
+$\lambda\cdot\alpha+(1-\lambda)\cdot\beta~,$
 subject to the linear constraints that define polytope $\mathcal{P}_{n,K,\,\tilde{F}}(\alpha,\beta)$. The optimal objective value $\mathrm{LP}^{*}(\lambda)$ then implies a hardness result of the form:
@@
 For details, see our code repository.
-Appendix A Missing Proofs for MaxProb Consistency
-A.1 Monotonicity of Consistency
+[^10] https://github.com/billyldc/optimal-stopping-with-a-predicted-prior
+Acknowledgement
+We are grateful for the insightful discussions with Yiding Feng, Enze Sun, and Xiaowei Wu.
+We also wish to thank the participants of the Dagstuhl workshop 2025 “Online Algorithms Beyond Competitive Analysis” and the “Trends in Approximation and Online Algorithms (TAO)” workshop at FOCS 2025 for their valuable feedback.
+Appendix A Derivation of Optimal Robust Threshold Functions
+In this section, we prove Lemma 8.
+First, we derive the indifference (Bellman) equations for the bi-criteria algorithm under the MaxExp and MaxProb objectives.
+Then, we show that, for each objective, these equations uniquely determine the optimal threshold functions $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ when the predicted prior coincides with the true prior $F$.
+We first treat the MaxExp objective and then turn to MaxProb.
+A.1 MaxExp: Optimal Threshold
+Lemma 21 (Optimal threshold function for MaxExp).
+For any $n\in\mathbb{N}$ and any prior $F$, the optimal threshold function $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ that maximizes the consistency of $\mathcal{A}(F,\theta)$ among all $\theta$’s under MaxExp is uniquely determined at each time $t$ by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$, where $x^{*}(t)$ is the unique solution of
+$\phi(t,x)~=~0,$ (17)
+with
+$\phi(t,x)~\coloneqq~\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]\left(1-\left(\frac{F(x)}{1-t+tF(x)}\right)^{n-1}\right)-x.$
+Moreover, $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$ is non-increasing in $t$ on $[0,1]$.
+Proof.
+Fix $t\in[0,1]$ and suppose the thresholds $\theta(s)$ have already been fixed for all $s\in(t,1]$.
+Let $x$ denote the value that arrives at time $t$, and assume $x$ is a record (else the algorithm cannot accept at $t$).
+By definition, the algorithm $\mathcal{A}(F,\theta_{n}^{\scriptscriptstyle\mathrm{exp}})$ accepts this value $x$ at time $t$ if and only if $F(x)>\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$.
+By optimality, the threshold chosen at time $t$ must make the expected rewards from accepting and rejecting this value $x$ equal.
+Accepting. Accepting yields a reward $x$.
+Rejecting. If the algorithm rejects $x$, then by the non-increasing threshold property, it must accept the first later value $y\geq x$. Thus, the continuation reward is
+$\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]\;\cdot\;\operatorname{\mathbf{Pr}}\!\big[\exists\,y\geq x\text{ after }t\mid x\text{ is a record at }t\big].$
+To compute the existence probability, note that all other n1n-1 samples lie below $x$ with probability $F(x)^{n-1}$, and each sample is either below $x$ or arrives after $t$ with probability 1t+tF(x)1-t+tF(x).
+Conditioning on the latter event gives
+$\operatorname{\mathbf{Pr}}\!\big[\exists\,y\geq x\text{ after }t\mid x\text{ is a record at }t\big]=1-\frac{F(x)^{n-1}}{(1-t+tF(x))^{\,n-1}}.$
+This conditioning introduces the factor $(1-t+tF(x))^{n-1}$, because each of the other n1n-1 values must either arrive after time $t$ or fall below $x$.
+Equating the accept and reject rewards yields exactly the equation $\phi(t,x)=0$ in Equation 17.
+By Lemma 22, for any continuous prior $F$ with full support on $0,\infty)$, any $n\in\mathbb{N}$, and any $t\in[0,1]$, the equation $\phi(t,x)=0$ has a unique solution $x^{*}(t)\in[0,\infty)$.
+Thus, defining $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)\coloneqq F(x^{*}(t))$ yields a unique optimal threshold at each time $t$.
+Finally, by Lemma 23, $x^{*}(t)$ is non-increasing in $t$ on $[0,1]$. Since $F$ is non-decreasing, the same holds for $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$.
+This completes the proof. ∎
+We now establish existence, uniqueness, and monotonicity of the solution $x^{*}(t)$ to Equation 17.
+For the proofs it will be convenient to introduce some auxiliary notation.
+For $u\in[0,1)$ and $t\in[0,1)$, define
+$g(u,t)\coloneqq\frac{1}{1-u}\left(1-\left(\frac{u}{1-t+tu}\right)^{n-1}\right),$
+and for any $(t,x)$ with $t\in[0,1)$ and $x\geq 0$ define
+$h(t,x)\coloneqq\frac{x}{g(F(x),t)}-\int_{x}^{\infty}y\,\,\mathrm{d}F(y).$
+Using
+$\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]=\frac{\int_{x}^{\infty}y\,\,\mathrm{d}F(y)}{1-F(x)},$
+one checks that for each fixed $t\in[0,1)$ and $x\geq 0$,
+$\phi(t,x)=0\quad\Longleftrightarrow\quad h(t,x)=0.$ (18)
+We first prove the existence and uniqueness of $x^{*}(t)$ for each $t$, and then show that $x^{*}(t)$ is non-increasing in $t$.
 Lemma 22.
-For any fixed $\lambda_{1},\lambda_{2}\in[0,1]$, the consistency $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is non-increasing in $n$.
+For any continuous prior $F$ with full support $0,\infty)$, any $n\in\mathbb{N}$, and any $t\in[0,1]$, Equation 17 has a unique solution $x^{*}(t)\in[0,\infty)$.
 Proof.
-We show that for any $n$, it holds:
-$\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\leq~\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~.$
-Consider a relaxed version of the problem with $n+1$ values, where the adversary tells the algorithm at the beginning the information about the smallest realized value, including its arrival time $\tau$ and its value $x_{\tau}$.
-We consider two policies.
-In the first policy, we ignore the additional information and directly apply the algorithm parameterized by $\theta^{*}_{n+1}$.
-By definition, the consistency of this policy is $\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
-We now describe the second and optimal policy.
-Conditioned on $\tau$ and $x_{\tau}$, the remaining $n$ values are i.i.d.
-from $F$ restricted to $x_{\tau},+\infty)$, and their arrival times are i.i.d.
-uniform on $[0,1]$.
-Thus, by the optimality of $\theta^{*}_{n}$, any algorithm that rejects all values during $0,\lambda_{1}]$, accepts best-so-far values during $\lambda_{1},\lambda_{2}]$ based on a threshold function, and accepts best-so-far values during $\lambda_{2},1]$ has consistency at most $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
-Since the second strategy is optimal, we get the stated inequality. ∎
-A.2 Asymptotic Characterization of $\theta_{n}^{*}$
+When $t=1$, the factor
+$1-\left(\frac{F(x)}{1-t+tF(x)}\right)^{n-1}$
+equals 0 for all $x\geq 0$, so $\phi(1,x)=-x$ has a unique zero $x^{*}(1)=0$. We focus on the case $0\leq t<1$ below.
+For fixed $t\in[0,1)$, the equivalence (18) implies that it suffices to show that $h(t,x)$ is strictly increasing in $x$ on $0,\infty)$ and crosses zero exactly once.
+First, $g(F(0),t)=g(0,t)=1$, and by L’Hôpital’s rule
+$\lim_{x\to\infty}g(F(x),t)=\lim_{u\to 1}g(u,t)=(n-1)(1-t).$
+Hence
+$h(t,0)$ $=-\int_{0}^{\infty}y\,\,\mathrm{d}F(y)\leq 0,$
+$\lim_{x\to\infty}h(t,x)$ $=\lim_{x\to\infty}\frac{x}{(n-1)(1-t)}=\infty,$
+so, by the Intermediate Value Theorem, $h(t,\cdot)$ has at least one root in $0,\infty)$.
+We now show that $h(t,x)$ has at most one root in $0,\infty)$. It suffices to prove that $\partial_{x}h(t,x)>0$ for any $x\geq 0$. Differentiating in $x$, we obtain
+$\partial_{x}h(t,x)=\frac{1}{g(F(x),t)}+\frac{x\,F^{\prime}(x)\big(g(F(x),t)^{2}-\partial_{u}g(F(x),t)\big)}{g(F(x),t)^{2}},$
+where $F^{\prime}(x)$ denotes the density of $F$ (which exists almost everywhere). Since $g(u,t)>0$ for any $u\in[0,1)$, it remains to show that for any $u\in[0,1)$,
+$g(u,t)^{2}-\partial_{u}g(u,t)\geq 0.$
+Differentiating in $u$,
+$g(u,t)^{2}-\partial_{u}g(u,t)=\frac{u^{n-2}\left[u^{n}-u(1-t+tu)^{n-1}+(n-1)(1-t)(1-u)(1-t+tu)^{n-2}\right]}{(1-u)^{2}(1-t+tu)^{2n-2}}.$
+Let $z(u,t)$ denote the bracketed term in the numerator. Then $z(u,1)=0$ for any $u\in[0,1)$, and
+$\partial_{t}z(u,t)=-(n-1)(1-u)^{2}(1-t+tu)^{n-3}\left[(1-t+tu)+(1-t)(n-2)\right]\leq 0.$
+Thus $z(u,t)\geq z(u,1)=0$ for any $u,t\in[0,1)$, which implies $g(u,t)^{2}-\partial_{u}g(u,t)\geq 0$ and hence $\partial_{x}h(t,x)>0$ for all $x\geq 0$.
+Therefore, for each fixed $t\in[0,1)$, the function $h(t,\cdot)$ is strictly increasing on $0,\infty)$ and crosses zero exactly once.
+By (18), the equation $h(t,x)=0$ has a unique solution $x^{*}(t)\in[0,\infty)$ for each fixed $t\in[0,1)$; and together with the case $t=1$, this proves the claim and the lemma.
+∎
 Lemma 23.
-For any fixed $s\in[0,1)$, we have
-$\theta^{*}_{n}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
-where $\gamma\approx 0.80435$ satisfies
-$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$
+For any continuous prior $F$ with full support $0,\infty)$ and any $n\in\mathbb{N}$, the unique solution $x^{*}(t)$ of Equation 17 is non-increasing in $t$ on $[0,1]$.
 Proof.
-Recall Equation 5:
-$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta^{*}_{n}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$
-Fix any $s$, substitute $x_{n}=\frac{1}{\theta^{*}_{n}(s)}-1$, the left-hand side becomes
-$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
-Expanding the left-hand side with the binomial theorem, we have
-$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k!}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{(n-1)!}x_{n}^{n-1}=1~.$
-Now, we analyze the asymptotic behavior of $x_{n}$.
-Obviously, we have $x_{n}=\mathcal{O}(\tfrac{1}{n-1})$; otherwise, the left-hand side either diverges to infinity or converges to 0, thus cannot equal $1$.
-Now, we find the constant such that $(n-1)(1-s)x_{n}\to c$ for some constant $c$ as $n\to\infty$. Note that such a constant $c$ shall satisfy
-$c+\frac{n-2}{n-1}\cdot\frac{c^{2}}{2!2}+\cdots+\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\cdot\frac{c^{k}}{k!k}+\cdots=1.$
-Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that $c$ satisfies
-$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
-which defines the constant $\gamma\approx 0.80435$ in the lemma.
-We can prove the convergence more rigorously by showing that, for any $c\in(0,1)$, the difference between the ideal and the actual sum tends to 0 as $n\to\infty$.
-Consider the difference:
-$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\left(1-\left(1-\frac{1}{n-1}\right)\left(1-\frac{2}{n-1}\right)\cdots\left(1-\frac{k-1}{n-1}\right)\right)~.$
-Observe that $1\leq(1-x_{1})(1-x_{2})\cdots(1-x_{m})+x_{1}+x_{2}+\cdots+x_{m}$ for any $x_{i}\in[0,1]$.
-Indeed, the partial derivative shows that the right-hand side is increasing in each $x_{i}$, thus the minimum is attained at $x_{i}=0$ for all $i$, which is $1$.
-It follows that the difference is bounded by
-$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{1+2+\cdots+(k-1)}{n-1}=\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{k(k-1)}{2(n-1)}\leq\frac{1}{2(n-1)}\sum_{k=1}^{\infty}\frac{1}{(k-1)!}~.$
-Using the fact that $\sum_{k=1}^{\infty}\frac{1}{(k-1)!}=\mathrm{e}$, the bound is $\mathcal{O}(\tfrac{1}{n})$ and therefore tends to 0 as $n\to\infty$.
-Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$.
-This completes the proof by $\theta^{*}_{n}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$.
+With $g(u,t)$ and $h(t,x)$ defined above, Lemma 22 implies that for each fixed $t\in[0,1)$ the function $h(t,\cdot)$ is strictly increasing on $0,\infty)$ and has a unique zero at $x=x^{*}(t)$.
+Fix $x\geq 0$ and let $u=F(x)$. For $u\in[0,1)$ and $t\in[0,1)$,
+$\partial_{t}g(u,t)=-\frac{(n-1)u^{n-1}}{(1-t+tu)^{n}}\leq 0,$
+so $g(F(x),t)$ is non-increasing in $t$, $1/g(F(x),t)$ is non-decreasing, and hence
+$\partial_{t}h(t,x)=x\,\partial_{t}\left[\frac{1}{g(F(x),t)}\right]\geq 0.$
+Thus, for any $0\leq t_{1}<t_{2}\leq 1$ and any $x\geq 0$,
+$h(t_{1},x)\leq h(t_{2},x).$
+Fix $0\leq t_{1}<t_{2}\leq 1$.
+Since $x^{*}(t_{2})$ is the unique zero of $h(t_{2},\cdot)$, we have $h(t_{2},x^{*}(t_{2}))=0$, and the inequality above gives $h(t_{1},x^{*}(t_{2}))\leq 0$.
+Because $h(t_{1},\cdot)$ is strictly increasing on $0,\infty)$ and has a unique zero at $x^{*}(t_{1})$, this implies $x^{*}(t_{1})\geq x^{*}(t_{2})$.
+Hence $x^{*}(t)$ is non-increasing in $t$ on $[0,1]$. ∎
+A.2 MaxProb: Optimal Threshold
+Lemma 24 (Optimal threshold function for MaxProb).
+For any $n\in\mathbb{N}$ and any prior $F$, the optimal threshold function $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ that maximizes the consistency of $\mathcal{A}(F,\theta_{n}^{\scriptscriptstyle\mathrm{prob}})$ under MaxProb is uniquely determined at each time $t$ by the unique solution $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)\in[0,1]$ of
+$\psi(t,q)~=~0,$ (19)
+where
+$\psi(t,q)~\coloneqq~\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s-q^{n-1}.$
+Moreover, $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing in $t$ on $[0,1]$.
+Note that the form of Equation (19) coincides with the equation defining the threshold function used by Example 2.
+Proof.
+Fix $t\in[0,1]$ and suppose the thresholds $\theta(s)$ have already been fixed for all $s\in(t,1]$.
+Let $x$ denote the value that arrives at time $t$, and assume $x$ is a record (otherwise the algorithm cannot accept at $t$).
+By definition, the algorithm $\mathcal{A}(F,\theta)$ accepts this value $x$ at time $t$ if and only if $F(x)>\theta(t)$.
+By optimality, the threshold chosen at time $t$ must make the success probabilities from accepting and rejecting this value $x$ equal.
+Accepting.
+If the algorithm accepts a record $x$ at time $t$, the success event is that $x$ is the largest among all $n$ values.
+Conditional on $x$ being a record at time $t$, this occurs with probability
+[…]
+Rejecting.
+If instead the algorithm rejects such an $x$, then by the non-increasing threshold assumption it will accept the next value larger than $x$, which is necessarily the next record that exceeds the threshold.
+Given that there are in total $k$ values larger than $x$, the probability that the algorithm eventually picks the largest among all values is $1/k$, since the arrival order is uniform at random and the algorithm accepts the first exceedance over $x$.
+The probability that there are exactly $k$ values larger than $x$ and arriving after time $t$, conditional on $x$ being a record at time $t$, is
+$\frac{\binom{n-1}{k}\big((1-t)(1-F(x))\big)^{k}F(x)^{n-1-k}}{(1-t+tF(x))^{n-1}}.$
+Equating the success probability from accepting and rejecting $x$ yields the backward-induction equation
+$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-F(x))\big)^{k}F(x)^{n-1-k}~=~F(x)^{n-1}.$ (20)
+For the quantile threshold $q=F(x)$, this becomes
+$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-q)\big)^{k}q^{n-1-k}~=~q^{n-1}.$
+Using the identity $\frac{1}{k}=\int_{0}^{1}y^{k-1}\,\,\mathrm{d}y$, the left-hand side can be written as
+$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-q)\big)^{k}q^{n-1-k}$ $=\sum_{k=1}^{n-1}\binom{n-1}{k}q^{n-1-k}\int_{0}^{(1-t)(1-q)}y^{k-1}\,\,\mathrm{d}y$
+$=\int_{0}^{(1-t)(1-q)}\frac{1}{y}\sum_{k=1}^{n-1}\binom{n-1}{k}q^{n-1-k}y^{k}\,\,\mathrm{d}y$
+$=\int_{0}^{(1-t)(1-q)}\frac{(y+q)^{n-1}-q^{n-1}}{y}\,\,\mathrm{d}y$
+$=\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s,$
+where the last equality follows from the change of variables y=(1s)(1q)y=(1-s)(1-q).
+Under this substitution, $y=0$ corresponds to $s=1$ and y=(1t)(1q)y=(1-t)(1-q) corresponds to $s=t$, which reverses the limits.
+Thus, Equation 20 is equivalent to
+$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s~=~q^{n-1},$
+which is precisely the equation $\psi(t,q)=0$ in Equation 19.
+We now prove existence and uniqueness of the solution of $\psi(t,q)=0$ for each fixed $t$, and then establish the monotonicity of $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$.
+For each fixed $t\in[0,1)$, the function $\psi(t,q)$ is continuous on $q\in[0,1]$. At $q=0$, the integrand equals $(1-s)^{n-2}$, hence
+$\psi(t,0)=\int_{t}^{1}(1-s)^{n-2}\,\,\mathrm{d}s>0.$
+At $q=1$, the integrand is identically zero, so
+$\psi(t,1)=0-1^{n-1}=-1<0.$
+Thus, for each $t\in[0,1)$ the equation $\psi(t,q)=0$ has at least one root $q\in(0,1)$.
+For uniqueness, observe first that when $t=1$, the integrand in $\psi(1,q)$ is zero for all $q\in[0,1]$, so $\psi(1,q)=-q^{n-1}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(1)=0$ is the unique root.
+For $t<1$, define
+$\tilde{\psi}(t,q)\coloneqq q^{-(n-1)}\psi(t,q)=\int_{t}^{1}\frac{\left(\frac{1-s+sq}{q}\right)^{n-1}-1}{1-s}\,\,\mathrm{d}s-1,\qquad q\in(0,1).$
+Differentiating with respect to $q$ gives
+$\partial_{q}\tilde{\psi}(t,q)=-\frac{n-1}{q^{n}}\int_{t}^{1}(1-s+sq)^{n-2}\,\,\mathrm{d}s<0,$
+so for each fixed $t<1$ the function $\tilde{\psi}(t,q)$ is strictly decreasing in $q$ on $(0,1)$.
+Thus, for each $t\in[0,1)$ the equation $\tilde{\psi}(t,q)=0$ has a unique solution $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)\in(0,1)$.
+Because $\tilde{\psi}(t,q)=0$ is equivalent to $\psi(t,q)=0$ for $q>0$, the solution of $\psi(t,q)=0$ is also unique.
+Finally, it remains to show that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing in $t$.
+Fix $0\leq t_{1}<t_{2}<1$ and suppose, for contradiction, that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$.
+Since $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{i})$ is the unique zero of $\tilde{\psi}(t_{i},q)$, we have
+$\tilde{\psi}(t_{i},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{i}))=0,\qquad i=1,2.$
+For fixed $t_{1}$, the function $q\mapsto\tilde{\psi}(t_{1},q)$ is strictly decreasing on $(0,1)$ and has a unique zero at $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$.
+The inequality $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ therefore implies
+$\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))<\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1}))=0.$ (21)
+On the other hand, for fixed $q=\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})$ the function $\tilde{\psi}(t,q)$ is strictly decreasing on $[0,1)$. Since $t_{1}<t_{2}$, this yields
+$\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))>\tilde{\psi}(t_{2},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))=0.$ (22)
+The inequalities (21) and (22) contradict each other.
+Hence the assumption $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ is impossible, and we must have $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})\leq\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ for all $0\leq t_{1}<t_{2}<1$.
+Thus $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing on $[0,1)$, and together with $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(1)=0$ this completes the proof.
 ∎
@@
 $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\theta\big(\max\{s,z\}\big)^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall z\in[\lambda_{1},\lambda_{2}]~.$ (24)
-We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in[\lambda_{2},1]$.
+We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in(\lambda_{2},1]$.
 Define the left-hand side of Equation 24 as $L(z)$ and simplify:
@@
 $~=$ $~z\int_{z}^{1}\frac{\theta(z)^{t}}{t}\,\mathrm{d}t+\int_{z}^{\lambda_{2}}\int_{s}^{1}\frac{\theta(s)^{t}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
-Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $\lambda_{2},1]$.
+Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $(\lambda_{2},1]$.
 We approximate $\theta$ by a left‑continuous non‑increasing step function computed backward from $\lambda_{2}$ to $\lambda_{1}$ so that Equation 24 holds for all $z\in[\lambda_{1},\lambda_{2}]$.
@@
 Thus, there exists a $\theta_{m}^{*}>0$ such that the equality holds.
-[^9] During experiment, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
+[^11] During experiments, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
 Backward recursion.
@@
 Proof.
-By unifying intervals that $\theta_{i}$ exceeds $1$, as Lemma 22, we only need to check that Equation 21 holds for $\min\{\theta_{i},1\}$, where $i$ is the last $\theta_{i}$ exceeding $1$.
-This instantly stands by the non-increasing of the left-hand side.
+By merging intervals where $\theta_{i}$ exceeds $1$, and by Lemma 26, we only need to check that Equation 27 holds for $\min\{\theta_{i},1\}$, where $i$ is the last index such that $\theta_{i}>1$.
+This follows immediately from the non-increasing property of the left-hand side.
 ∎
@@
 Corollary 29.
-If an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 17 and Equation 18.
+If there exists an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 23 and Equation 24.
 In our implementation, we fix several pairs of values $(\alpha,\beta)$ and numerically verify that for each pair, the constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$.
@@
 For details, see the implementation of the backward recursion as well as the solved $\theta_{i}^{*}$’s in our code repository.
-Figure 3 illustrates the case when $\beta=0.01$ and $m=300$ yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$.
+[^12] https://github.com/billyldc/optimal-stopping-with-a-predicted-prior
+Figure 3 illustrates the case where $\beta=0.01$ and $m=300$, which yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$.
 This implies that $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ with $\theta^{*}$ being the step function with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{300}$ is an $\alpha\approx 0.6908$-consistent and $\beta=0.01$-robust algorithm.
 Figure 3: Example solution of $\theta$ when $\beta=0.01$, $m=300$, and $\alpha\approx 0.6908$.
-Appendix C Missing Proofs for MaxProb Hardness
-C.1 Minor-Oblivious Algorithms: Proof of Proposition 17
-The key observation is that, when the current value is best-so-far and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability.
+Appendix C Missing Proofs for MaxProb Consistency
+Lemma 30.
+For any fixed $s\in[0,1)$, we have
+$\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
+where $\gamma\approx 0.80435$ satisfies
+$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$
+Proof.
+Recall Equation 28:
+$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$ (28)
+Fix any $s$, substitute $x_{n}=\frac{1}{\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)}-1$, the left-hand side becomes
+$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
+Expanding the left-hand side with the binomial theorem, we have
+$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{n-1}x_{n}^{n-1}=1~.$
+Now, we analyze the asymptotic behavior of $x_{n}$.
+Obviously, we have $x_{n}=\mathcal{O}(\tfrac{1}{n-1})$; otherwise, the left-hand side either diverges to infinity or converges to 0, thus cannot equal $1$.
+Now, we find the constant such that $(n-1)(1-s)x_{n}\to c$ for some constant $c$ as $n\to\infty$. Note that such a constant $c$ shall satisfy
+$c+\frac{n-2}{n-1}\cdot\frac{c^{2}}{2!2}+\cdots+\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\cdot\frac{c^{k}}{k!k}+\cdots=1.$
+Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that the value $c$ satisfies
+$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
+which defines the constant $\gamma\approx 0.80435$ in the lemma.
+We can prove the convergence more rigorously by showing that, for any $c\in(0,1)$, the difference between the ideal and the actual sum tends to 0 as $n\to\infty$.
+Consider the difference:
+$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\left(1-\left(1-\frac{1}{n-1}\right)\left(1-\frac{2}{n-1}\right)\cdots\left(1-\frac{k-1}{n-1}\right)\right)~.$
+Observe that $1\leq(1-x_{1})(1-x_{2})\cdots(1-x_{m})+x_{1}+x_{2}+\cdots+x_{m}$ for any $x_{i}\in[0,1]$.
+Indeed, the partial derivative shows that the right-hand side is increasing in each $x_{i}$, thus the minimum is attained at $x_{i}=0$ for all $i$, which is $1$.
+It follows that the difference is bounded by
+$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{1+2+\cdots+(k-1)}{n-1}=\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{k(k-1)}{2(n-1)}\leq\frac{1}{2(n-1)}\sum_{k=1}^{\infty}\frac{1}{(k-1)!}~.$
+Using the fact that $\sum_{k=1}^{\infty}\frac{1}{(k-1)!}=\mathrm{e}$, the bound is $\mathcal{O}(\tfrac{1}{n})$ and therefore tends to 0 as $n\to\infty$.
+Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$.
+This completes the proof by $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$.
+∎
+Appendix D Missing Proofs for MaxProb Hardness
+D.1 Minor-Oblivious Algorithms: Proof of Proposition 19
+The key observation is that, when the current value is a record and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability.
 This is formalized in the following lemma.
@@
 $r^{\prime}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t,\ x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~.$
-That is, the probability of accepting a best-so-far value $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is best-so-far; and 3) $\mathcal{A}$ rejects the first t1t-1 values.
+That is, the probability of accepting a record $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is the record; and 3) $\mathcal{A}$ rejects the first t1t-1 values.
 By Lemma 31, we may define this conditional probability with respect to prior $F_{K}$ without loss of generality.
@@
 The proof of Equation 30 for $t+1$ is almost verbatim.
-C.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 18
+D.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 20
 From Rejection Probabilities to Acceptance Probabilities.